### v1

# From Coarse to Fine-Grained Emotion Annotation: An Immediate Recall Paradigm with Validation through Physiological Evidence and Recognition Performance

Abstract
Traditional video-induced emotion physiological datasets often use whole-trial annotation, assigning a single emotion label to all data collected during an entire trial. This coarse-grained annotation approach misaligns with the dynamic and temporally localized nature of emotional responses as they unfold with video narratives, introducing label noise that limits emotion recognition algorithm evaluation and performance. To solve the label noise problem caused by coarse-grained annotation, we propose a fine-grained annotation method through an immediate recall paradigm. This paradigm integrates an immediate video replay phase after the initial stimulus viewing, allowing participants to precisely mark the onset timestamp, emotion label, and intensity based on their immediate recall. We validate this paradigm through physiological evidence and recognition performance. Physiological validation of multimodal signals within participant-marked windows revealed rhythm-specific EEG patterns and arousal-dependent GSR responses—with SCRs appearing in 91% of high-arousal versus 6% of low-arousal emotion windows. These objective physiological data changes strongly aligned with subjective annotations, confirming annotation precision. For recognition performance, classification experiments showed that models trained on fine-grained annotations achieved 9.7% higher accuracy than traditional whole-trial labeling, despite using less data. This work not only addresses label noise through fine-grained annotation but also demonstrates that annotation precision outweighs data scale in determining emotion recognition performance.
Index Terms:
Emotion Annotation, Label Noise, Multimodal Emotion Dataset, Physiological Validation, Emotion Recognition.
I Introduction
Emotion is a subjective experience elicited by external stimuli [1], typically accompanied by physiological responses and behavioral tendencies. Emotions play a crucial role in individual survival, social interaction [2], and decision-making processes [3]. Two primary theoretical frameworks describe emotions: discrete and dimensional models. The discrete model, exemplified by Ekman’s seminal work [4], identifies six basic emotions—happiness, sadness, fear, anger, surprise, and disgust—each linked to specific neural circuits. For instance, fear demonstrates strong associations with amygdala activation. In contrast, the dimensional model, represented by Russell’s Circumplex Model [5], characterizes emotions along two dimensions: valence and arousal. These complementary frameworks establish a robust theoretical foundation for emotion research.
To explore neural mechanisms and develop recognition methods for emotions, researchers have designed various experimental paradigms to simulate and capture human emotional responses. These experiments typically employ music [6], images [7], or videos [8] as stimuli while collecting participants’ physiological and behavioral signals. Among these modalities, video stimuli demonstrate superior performance in eliciting complex emotions due to their dynamic nature and multisensory input, thereby offering higher ecological validity compared to static images or single-modality audio stimuli [9]. Data collection encompasses both physiological signals (e.g., electroencephalogram (EEG), electrocardiogram (ECG)) [10, 11] and non-physiological signals (e.g., facial expressions) [12], providing comprehensive documentation of emotional manifestations. Physiological signals offer high temporal resolution and objectivity, enabling precise capture of neural and physiological changes associated with emotions [13]. Non-physiological signals provide behavioral cues that complement physiological measures, further deepening our multidimensional understanding of emotional mechanisms [14].
Despite continuous developments in experimental paradigms for emotion research, existing methods face significant challenges in emotion annotation accuracy. Datasets such as SEED [15] and DEAP [16] employ coarse-grained annotation strategies, assigning single emotion labels to entire multi-minute trials. While this approach reduces annotation time, it overlooks the temporal instability and discrete nature of emotions, resulting in substantial label noise. To address this limitation, Zhang et al. [17] proposed an emotion localization task for the SEED dataset. They segmented each video stimulus into multiple intervals and recruited 20 annotators to evaluate whether each segment triggered the target emotion. Segments receiving majority agreement were labeled as ground truth, thereby reducing noise by focusing on highly relevant emotional segments. Building upon this foundation [17], Zhang et al. [18] introduced TAS-Net, an unsupervised time-aware sampling network based on deep reinforcement learning. This approach automatically detects key emotional segments in EEG signals while reducing labeling noise from traditional fixed-label methods. The method’s accuracy was validated by comparing detected segments with manually re-annotated segments. Similarly, Jiang et al. proposed REMONET [19], which mitigates sentiment label noise through a multi-regularization co-learning framework that coordinates different self-supervised data views to prevent model overfitting to noisy labels. However, these post-hoc approaches—whether through manual re-annotation or deep learning techniques—do not fundamentally resolve the inherent ambiguity in mapping physiological data to subjective emotion labels [20].
In recent efforts to address annotation granularity, researchers have developed multimodal datasets such as SEED-VII [21] and MGEED [22], aiming to achieve finer-grained emotional labeling. SEED-VII represents a significant advancement by incorporating both trial-by-trial intensity ratings and post-session continuous annotation. Specifically, participants provide immediate intensity feedback after each trial, then retrospectively assign frame-level emotional intensities while reviewing all videos at the session’s conclusion. This dual-annotation approach attempts to capture both immediate impressions and detailed temporal dynamics. Similarly, MGEED employs a comprehensive frame-by-frame retrospective protocol, where participants meticulously annotate their emotional states for each video frame after viewing all experimental stimuli. While these methodologies achieve fine-grained temporal annotation, both datasets fundamentally rely on delayed recall mechanisms. This “watch first, annotate later” paradigm, despite its temporal granularity, introduces significant cognitive load [23], as participants must accurately remember and differentiate emotional experiences across multiple stimuli viewed over extended periods. Moreover, inherent memory bias [24] in delayed reporting may lead to temporal displacement of emotional events or conflation of distinct emotional episodes, potentially compromising annotation reliability.
To achieve fine-grained emotion annotation while mitigating memory bias, we propose an immediate recall annotation paradigm and establish the Fine-grained Immediate Recall-based Multimodal Emotion Dataset (FIRMED). This approach introduces three key methodological advancements:
1) Precise Temporal Localization via Immediate Recall: To achieve precise temporal localization of emotional events within each trial, our paradigm incorporates an immediate replay phase following stimulus presentation. During post-stimulus replay, participants retrospectively identify and mark the discrete time point $t_{\text{event}}$ at which they recall first experiencing the target emotion during initial viewing. A fixed-duration window (4 seconds) centered at $t_{\text{event}}$ is then extracted for analysis, associating data within this window with the reported emotion label and intensity. This recall-based discrete annotation during immediate replay aims to overcome temporal ambiguity in whole-trial labeling while reducing memory decay biases associated with delayed annotation protocols.
2) Expanded Multimodal Data for Enhanced Reliability: To improve annotation reliability and objectivity, we expanded the range of collected physiological signals to include EEG, ECG, galvanic skin response (GSR), and photoplethysmogram (PPG) data. Additionally, we captured participants’ facial expressions via camera during stimulus viewing, providing behavioral data on emotional responses. This multimodal collection strategy enriches data variety and enables cross-validation of subjective annotations using various physiological data.
3) Ecologically Valid Stimuli: We extensively utilize short videos from social media platforms as stimuli, encompassing authentic scenes including daily life recordings and news reports. Compared to traditional movie clips or music videos, these stimuli more accurately reflect natural emotional experiences, offering enhanced ecological validity. Participant feedback regarding stimulus effectiveness further supports this choice.
To validate this paradigm’s effectiveness, we conducted systematic physiological and computational analyses. Physiologically, by analyzing multimodal signals within precisely annotated windows, we found strong evidence of both central [25] and autonomic nervous system engagement [26]. This included discovering emotion-specific spectral patterns in EEG signals and strong temporal coupling between discrete annotations and objective autonomic nervous system activation markers, exemplified by specific skin conductance response (SCR) patterns reliably occurring for high-arousal emotions [27]. Computationally, emotion classification experiments compared the effectiveness of the two labeling strategies. Results demonstrate that standard classifiers trained using precise labels from FIRMED significantly outperform identical models trained using traditional whole-trial labels applied to corresponding full trial segments from the same data, thereby quantifying the positive impact of reduced label noise on recognition performance.
II Related Work
II-A Evolution of Physiological Signal-Based Emotion Datasets
Video-based emotion elicitation has established itself as the predominant stimulus modality for inducing emotional responses in controlled experimental environments, driving the proliferation of multimodal datasets across the affective computing domain [28]. These datasets systematically collect multiple physiological signals to provide comprehensive documentation of participants’ emotional states [29].
Early and highly influential datasets, including DEAP [16] and the foundational SEED series [15, 30, 31], established the methodological foundation for physiological emotion analysis through rich signal acquisition protocols. These seminal works adopted a standardized experimental paradigm involving presentation of emotionally evocative stimuli—typically music videos or film clips—followed by collection of self-reported emotional assessments. However, these datasets universally employ whole-trial labeling, wherein single emotion labels are assigned to complete multi-minute stimulus segments.
This coarse-grained annotation strategy ignored the inherently dynamic nature of video stimuli. Emotional responses are characteristically triggered by specific narrative elements—including pivotal plot developments, climactic sequences, or sudden scene transitions—rather than being uniformly sustained throughout extended video segments [32]. Consequently, this labeling granularity introduces label noise [19], as temporal segments that fail to elicit the target emotion become erroneously associated with that emotional state.
In direct response to these temporal granularity limitations, recent dataset development efforts have pursued increasingly sophisticated annotation methodologies. The SEED-VII dataset [21] introduced a delayed continuous rating paradigm representing a significant methodological advancement. In this approach, participants retrospectively review all trial videos upon session completion and provide continuous intensity annotations. Similarly, the MGEED database [22] implemented a delayed frame-by-frame self-assessment protocol, requiring participants to annotate their emotional states retrospectively following complete viewing session.
To systematically characterize these dataset developments, Table I provides a comprehensive comparison of representative emotion datasets. The analysis reveals a clear evolutionary trajectory from coarse whole-trial labeling toward increasingly fine-grained temporal annotations.
TABLE I: Evolution of Multimodal Emotion Datasets: Labeling Granularity Paradigm Progression
Year Dataset Experimental Setup Signals Emotion Model Labeling Strategy
Coarse-grained annotation paradigm
2011 MAHNOB-HCI [8] N=27
20 video clips
EEG (32-ch), ECG, GSR, EOG, RSP, A-V 6 Discrete emotions + VAD Whole-trial
(Single label per trial)
2012 DEAP [16] N=32
40 music videos (60s)
Single session
EEG (32-ch), EOG, EMG, GSR, BVP, Temp., RSP, Face VAD Whole-trial
(Single label per trial)
2013 SEED [15] N=15
45 clips ($\sim$4 min)
3 sessions
EEG (62-ch), EOG Positive-Negative-Neutral Whole-trial
(Single label per trial)
2017 DREAMER [33] N=23
18 clips (60s)
Single session
EEG (14-ch), ECG VAD Whole-trial
(Single label per trial)
2018 AMIGOS [34] N=40
20 videos
(Mixed duration)
EEG (14-ch), ECG, GSR, RGB-D VAD Whole-trial
(Single label per trial)
2019 MPED [35] N=23
28 clips (2.5-5 min)
2 sessions
EEG (62-ch), ECG, RSP, GSR 7 Discrete emotions Whole-trial
(Single label per trial)
2021 SEED-V [31] N=20
45 clips (2-4 min)
3 sessions
EEG (62-ch), EOG 5 Discrete emotions + Neutral Whole-trial
(Single label per trial)
Fine-grained annotation paradigm
2023 MGEED [22] N=17
9 clips ($\sim$4 min)
Single session
EEG (14-ch), ECG, GSR, Optomyography 6 Discrete emotions + VAD Frame-level
(Delayed post-session)
2024 SEED-VII [21] N=20
80 clips (2-5 min)
4 sessions
EEG (62-ch), EOG 6 Discrete emotions + Neutral Continuous
(Delayed post-session)
Abbreviations: VAD = Valence-Arousal-Dominance; RSP = Respiration; BVP = Blood Volume Pulse; Temp. = Temperature; A-V = Audio-Visual
II-B Challenges of Emotion Annotation
The systematic adoption of whole-trial labeling across early multimodal emotion datasets [16, 15, 34, 33, 8, 35]. This annotation approach suffers from temporal imprecision: emotions constitute dynamic phenomena that evolve synchronously with video narratives [32], exhibiting rapid fluctuations and transitions rather than maintaining static states across extended durations. The assignment of uniform labels to multi-minute segments obscures emotional transitions and fluctuations [36], causing label noise to be structurally embedded in the dataset.
While recent datasets such as SEED-VII and MGEED have achieved significant improvements in labeling granularity through delayed recall paradigms. These approaches require participants to complete entire experimental sessions—encompassing numerous stimuli viewed over extended periods (typically 30 minutes to over 1 hour)—before retrospectively annotating their emotional states across all preceding trials. This delayed annotation strategy, despite achieving fine temporal granularity, it also increases the risk of memory bias, which may reduce the reliability of annotations. Cognitive psychology research [37] demonstrates that episodic memory for specific emotional events and their nuanced temporal characteristics becomes increasingly unreliable when recall is required after extended delays involving multiple distinct experiences.
The persistence of these limitations—temporal imprecision in whole-trial labeling and memory bias in delayed fine-grained annotation—underscores the urgent need for innovative annotation paradigms capable of achieving fine temporal granularity while maintaining annotation reliability.
III Experimental Setup
This Section details the experimental setup and methodology employed in this study during the data collection and preparation stages. The content is divided into four subsections: Section III-A describes the stimuli used for emotion elicitation, Section III-B details the participant demographics and recruitment information, Section III-C elaborates on the complete experimental procedure including the annotation process, and Section III-D explains the data preprocessing steps applied to the acquired multimodal signals.
III-A Stimuli
To enhance ecological validity and elicit emotional responses more representative of real-world experiences, we primarily selected short videos sourced from social media platforms (e.g., Douyin, Bilibili, YouTube) as stimuli, each lasting approximately one minute. These videos were often user-generated or edited by creators. The content primarily included daily life recordings and news reports, and the videos typically contained multiple segments potentially capable of evoking specific emotions within the short timeframe.
The stimulus selection involved a rigorous screening process. First, a large pool of candidate videos was initially collected and downloaded from various online platforms by two members of the research team. Subsequently, these candidates were strictly evaluated by a panel of three independent raters who are psychology graduate students trained in emotion recognition. The selection criterion was: only videos for which all three raters unanimously agreed could effectively elicit one specific target basic emotion (Happiness, Sadness, Anger, Fear, Disgust, or Surprise) were ultimately chosen. Videos failing to reach unanimous agreement on a single target emotion or eliciting mixed emotions were excluded during this phase.
Through this rigorous screening process, we obtained the final stimulus set containing 90 selected videos. This set was balanced across the six basic emotions and these 90 videos were evenly distributed across 3 experimental sessions. Each session included 30 trials, with each trial featuring one video specifically selected to elicit one of the six target basic emotions, resulting in 5 distinct videos per emotion category per session. This stimulus selection strategy, prioritizing ecological validity and confirming elicitation of specific target emotions via consensus rating, was corroborated by informal post-experimental feedback from participants; they generally reported that these real-world clips were more engaging and significantly more effective in eliciting their emotional resonance compared to more traditional laboratory stimuli such as movie clips or music videos.
III-B Participants
A total of twenty healthy participants (12 males, 8 females) were recruited for this study. Their ages ranged from 22 to 30 years, with a mean age of 24.7 years ± 2.1 years (Mean ± SD). All participants reported being right-handed, having normal or corrected-to-normal vision, and reported no history of neurological or psychiatric disorders. This study received ethical approval from the Northwestern Polytechnical University Ethics Committee. Before the experiment, the procedure and purpose of the study were introduced to the participants. Subsequently, by signing a consent agreement, participants agreed to the recording, sharing, and publication of their physiological signals and facial images for academic purposes. Participants were aware that their participation was voluntary and that they were free to withdraw at any time without providing any reason.
III-C Protocol
Participants were comfortably seated in a quiet shielded room, at a distance of approximately 70 cm from the display monitor. After providing informed consent and receiving detailed instructions about the task, multimodal physiological sensors including an EEG cap (acquiring 59 channel signals at 1000 Hz), ECG electrodes (acquiring 3 channel signals at 1000 Hz), GSR sensors (acquiring 2 channel signals at 100 Hz), and a PPG sensor (acquiring 1 signal at 50 Hz) were attached by trained researchers, connected to the Neuracle NeuroHUB wearable multimodal research platform, as shown in Fig. 1. Signal quality was checked after placement. The facial expression recording camera, a Logitech BRIO 500, was positioned on the top of the monitor used for stimulus viewing.
Figure 1: Neuracle NeuroHUB wearable acquisition device wearing method and EEG electrode positions.
The overall experimental procedure is illustrated in Fig. 2. The experiment consisted of three separate sessions conducted on different days, with at least a 24-hour interval between sessions. Each session comprised 30 trials, presenting 5 distinct videos for each of the 6 basic emotions. The presentation followed a fixed sequence of six emotion categories: Happiness $\rightarrow$ Surprise $\rightarrow$ Disgust $\rightarrow$ Anger $\rightarrow$ Fear $\rightarrow$ Sadness. This block was repeated five times within the session. For each repetition cycle, a unique video for the corresponding emotion category in the sequence was selected without replacement from the pool of five available videos for that category. This design ensured all 30 videos (5 per emotion) were presented exactly once per session while maintaining the fixed progression of emotion types.
Figure 2: Experimental paradigm for FIRMED data collection, showing the multi-session design, the procedure within a single trial (including baseline, video stimulus presentation, immediate replay with recall-based discrete annotation, and rest), the multimodal physiological and behavioral signals recorded, and the dimensions of emotion annotation (label, intensity, time).
Each trial consisted of four phases:
(1) Baseline: A fixation cross was displayed on the screen for 10 seconds. Participants were asked to relax and minimize movement. Physiological and behavioral data corresponding to this period were collected.
(2) Stimulus Presentation: One of the about 1-minute emotion-eliciting video clips was presented. Participants were instructed to watch the video naturally and keep their bodies as still as possible. Physiological and facial data corresponding to this period were collected.
(3) Immediate Replay and Annotation: Immediately following stimulus presentation, the video was replayed. Participants recalled their emotional experience from the initial viewing and indicated to the experimenter when they first experienced the target emotion. Due to GSR and PPG sensors on their fingers preventing direct interface control, the experimenter operated the annotation software based on participant instructions. Using the interface controls shown in Fig. 3, including an adjustable progress bar, the experimenter navigated to the precise timestamp identified by the participant. Upon participant confirmation, the experimenter paused the replay and recorded the timestamp. Participants then reported the emotion label and intensity level (Low, Medium, High), which the experimenter entered into the annotation panel, thereby linking the temporal, categorical, and intensity dimensions of the emotional response.
Figure 3: Emotion annotation software interface. Software functions include playing videos, adjusting video progress, recording video progress time, emotion category and emotion intensity.
(4) Rest: A 15-second rest period followed the all annotation confirmation before the next trial began.
III-D Data preprocessing
Prior to analysis, the continuously recorded raw multimodal signals underwent filtering to remove noise and artifacts by MNE-Python [38]. Specifically:
EEG data: The EEG signals were first downsampled to 250 Hz to reduce computational load while preserving relevant frequency information. The data were then band-pass filtered between 0.5 Hz and 70 Hz using a finite impulse response (FIR) filter, followed by a 50 Hz notch filter to remove power line noise. Independent Component Analysis (ICA) [39] was employed to identify and remove typical artifacts including eye blinks (EOG) and muscle activity (EMG). Bad channels were identified using a variance-based criterion: channels with variance exceeding 10 times the median variance across all channels were marked as ’bad’ and subsequently repaired using spherical spline interpolation.
ECG data: ECG signals were band-pass filtered between 0.5 Hz and 40 Hz using a 4th-order Butterworth filter to remove baseline wander and high-frequency noise.
GSR data: GSR signals were preprocessed primarily focusing on the tonic component. For this, a low-pass filter was applied with a cutoff frequency of 0.5 Hz.
PPG data: PPG signal were band-pass filtered between 0.5 Hz and 8 Hz using a 4th-order Butterworth filter to retain the primary pulse wave frequencies while removing baseline drift and noise.
Subsequently, analysis epochs were extracted based on the discrete event timestamps $t_{event}$ identified during the annotation phase. Acknowledging that $t_{event}$ represents a subjective recollection of emotion onset and thus likely contains temporal jitter (due to factors such as memory precision limitations, reaction time during annotation, and the inherent difficulty in pinpointing the exact onset of a subjective feeling), and considering that physiological responses typically unfold over a period rather than being instantaneous, we employed a fixed window approach for analysis. Specifically, for each annotated $t_{event}$, a 4-second window was extracted from the preprocessed multimodal data streams [40], encompassing 2 seconds before and 2 seconds after $t_{event}$ (i.e., [$t_{event}-2\,\text{s},t_{event}+2\,\text{s}$]). This window duration was selected to capture the most salient physiological changes surrounding the reported emotional event while accommodating the potential temporal variability [41].
IV Physiological Validation of the Annotation Paradigm
In Section III, we detailed the experimental procedure, the novel annotation paradigm, and the preprocessing methods for the multimodal data. This Section presents in-depth analyses of the multimodal physiological signals based on the processed data, the corresponding annotations, and the time windows defined around the annotated events, aiming to physiologically validate the annotation effectiveness. The analysis comprises two main parts: first, the results of the EEG signal analysis in Section IV-A; second, the results of the peripheral physiological signal analysis in Section IV-B; and finally, Section IV-C quantitatively evaluates the comparative results of our immediate annotation paradigm relative to a delayed annotation paradigm.
IV-A EEG Analysis Results
This subsection details the EEG analysis results aimed at identifying brain neural correlates associated with the precisely annotated emotional event markers $t_{event}$ derived from the FIRMED. We focused on event-related changes in Power Spectral Density (PSD), comparing activity within the emotion-event time windows to the pre-trial baseline period [42]. PSD was chosen as it is a well-established method for quantifying the distribution of power across different frequency bands in EEG signals, providing insights into the neural oscillatory changes that are widely recognized as correlates of various cognitive and emotional states. We calculated the PSD for each channel across standard frequency bands: $\delta$ (1-4 Hz), $\theta$ (4-8 Hz), $\alpha$ (8-13 Hz), $\beta$ (13-30 Hz) and $\gamma$ (30-45 Hz). For each participant and trial, preprocessed EEG data were extracted for the emotion-event time windows and the 10s baseline window. PSD for each channel within each window was calculated using Welch’s method.
For each participant and emotion category, the difference change in PSD from baseline to the event window is shown in (1). Subsequently, to assess the statistical significance of these changes at the group level, one-sample t-tests were performed at each electrode location to test whether the group-averaged $\Delta PSD$ values differed significantly from zero. To address the multiple comparisons problem inherent in simultaneous testing across 59 EEG channels, we employed cluster-based permutation testing [43]. This approach involves forming spatial clusters of electrodes showing consistent effects based on their adjacency relationships, followed by 1000 random permutations where the sign of each participant’s $\Delta PSD$ values is randomly flipped to simulate the null hypothesis of no systematic effect. Cluster-level statistics (sum of t-values within each cluster) from the observed data are then compared against the distribution of maximum cluster statistics obtained from the permutations to determine statistical significance. This non-parametric statistical method effectively controls the family-wise error rate by taking into account the spatial dependencies between channels, thereby offering greater sensitivity to detect genuine, spatially coherent neural effects. The findings described below highlight patterns observed at the group level based on these average effects.
$\Delta PSD=PSD_{event}-PSD_{baseline}$ (1)
Happiness: Group-level analysis of happiness events revealed widespread neural oscillatory changes characterized by multi-frequency engagement as shown in Fig. 4. The $\alpha$ band demonstrated the most extensive response, with 48 of 59 electrodes (81.4%) forming a single significant cluster ($p=0.004$, $d=-0.63$ to $0.03$). This widespread $\alpha$ desynchronization spanned nearly the entire cortical surface, reflecting global cortical disinhibition and enhanced neural readiness characteristic of positive emotional states [44]. The $\gamma$ band showed remarkable synchronization, with 44 electrodes (74.6%) displaying significant power increases within a coherent cluster ($p=0.001$, $d=0.00$ to $1.11$). This extensive $\gamma$ enhancement, particularly prominent in frontal, temporal, parietal, and occipital regions, supports the neural binding hypothesis and suggests widespread cross-regional integration during positive emotional experiences [45]. The $\beta$ band exhibited a selective activation pattern involving 27 electrodes (45.8%) organized into two distinct clusters ($p=0.018$ and $p=0.001$, $d=-0.46$ to $1.17$). This bi-cluster organization likely reflects coordinated activation of prefrontal-parietal cognitive control and temporo-occipital emotion regulation networks.
Figure 4: Group-level statistical T-maps for the happiness events. Channels belonging to statistically significant clusters are marked with white circles. Greyed-out topomaps indicate no significant clusters were found for that band.
In contrast, lower frequency bands showed markedly different response patterns. The $\delta$ frequency band demonstrated limited but focused significant changes affecting only 5 out of 59 channels, primarily localized to left temporal and parieto-occipital region. The $\theta$ frequency band did not exhibit any statistically significant clustered effects, indicating that this frequency range did not show spatially coherent changes during happy emotion processing in this study.
Surprise: Surprise events demonstrated more constrained cortical activation compared to other emotions, with effects primarily confined to mid-to-high frequency bands as shown in Fig. 5. The $\alpha$ band showed the most prominent response, with 32 of 59 electrodes (54.2%) forming two distinct clusters ($p=0.011$ and $p=0.001$, $d=-0.87$ to $0.04$). This $\alpha$ desynchronization, primarily involving central sensorimotor and posterior parietal regions, was substantially more limited than the near-global patterns observed for emotions like happinsess, suggesting a more focused cortical response. The $\beta$ band exhibited selective activation with only 11 electrodes (18.6%) organized into two significant clusters ($p=0.015$ and $p=0.001$, $d=-0.65$ to $0.23$), likely reflecting limited engagement of frontal and parietal cognitive control regions necessary for rapid reorientation. The $\gamma$ band showed minimal response, with merely 5 electrodes (8.5%) forming a single cluster ($p=0.018$, $d=-0.53$ to $0.29$), indicating restricted high-frequency synchronization in specific attention-related areas.
Notably, both $\delta$ and $\theta$ bands showed no statistically significant clustered effects, clearly distinguishing surprise from emotions that engage low-frequency oscillations. This absence of low-frequency involvement, combined with moderate $\alpha$ desynchronization and limited high-frequency engagement, suggests that surprise represents a relatively simple and transient emotional state primarily characterized by rapid attentional reorientation rather than the deep neural integration observed in more complex emotional experiences [46].
Figure 5: Group-level statistical T-maps for the surprise events. Channels belonging to statistically significant clusters are marked with white circles. Greyed-out topomaps indicate no significant clusters were found for that band.
Disgust: Group-level analysis of Disgust events evealed a distinctive neurophysiological profile characterized by selective frequency-band modulations, most notably featuring $\theta$ band suppression as shown in Fig. 6. A unique finding was the significant power decrease in the $\theta$ band, localized to a cluster of 5 electrodes (8.5%) in the right posterior temporal-parietal-occipital region ($p=0.044$, $d=-1.04$ to $0.61$). This localized $\theta$ suppression, absent in all other analyzed emotions, may reflect specific inhibitory processes and attentional disengagement mechanisms associated with aversive stimulus processing. The $\alpha$ band demonstrated significant and widespread desynchronization, with 27 electrodes (45.8%) forming a single cluster ($p=0.025$). Notably, effect sizes were uniformly negative ($d=-0.77$ to $-0.07$), indicating consistent cortical activation across affected regions, primarily in central and posterior areas. This pattern suggests moderate but reliable cortical engagement during disgust processing, though less extensive than observed for high-arousal emotions. The $\beta$ band showed selective modulation involving 14 electrodes (23.7%) in a significant cluster ($p=0.003$, $d=-0.94$ to $0.70$), concentrated in parietal and parieto-occipital regions. The mixed positive and negative effect sizes suggest heterogeneous cortical responses, potentially reflecting the complex interplay between sensory rejection and motor preparation for avoidance behaviors.
No statistically significant clusters were identified in the $\delta$ or $\gamma$ bands after correction for multiple comparisons. This absence of low-frequency slow-wave activity and high-frequency binding distinguishes Disgust from emotions like Fear (characterized by $\delta$ enhancement) and Happy (marked by $\gamma$ synchronization). The distinct combination of $\theta$ suppression, moderate $\alpha$ desynchronization, and selective $\beta$ modulation constitutes the unique neurophysiological signature of disgust.
Figure 6: Group-level statistical T-maps for the disgust events. Channels belonging to statistically significant clusters are marked with white circles. Greyed-out topomaps indicate no significant clusters were found for that band.
Anger: Group-level analysis of anger events revealed the most restricted cortical activation pattern among all analyzed emotions, with significant modulations primarily confined to the $\alpha$ band as shown in Fig. 7. The $\alpha$ band showed moderate desynchronization, with 21 of 59 electrodes (35.6%) forming a single significant cluster ($p=0.004$, $d=-0.87$ to $-0.002$), concentrated over central, parietal, and occipital regions. This $\alpha$ suppression, while statistically significant, was markedly less extensive compared to emotions like happiness, suggesting a more localized cortical arousal pattern consistent with the focused nature of anger-related processing. The $\delta$ band demonstrated very focal activation with only 2 electrodes (3.4%) forming a significant cluster ($p=0.042$, $d=-0.52$ to $0.23$) in the left frontal area. Similarly, the $\beta$ band showed highly restricted activation involving merely 3 electrodes (5.1%) in a significant cluster ($p=0.035$, $d=-0.80$ to $0.24$), localized to right temporal regions.
Notably, both $\theta$ and $\gamma$ bands showed no statistically significant clustered effects after correction for multiple comparisons. The neural signature of anger in this paradigm was characterized by moderate $\alpha$ desynchronization with minimal involvement of other frequency bands, suggesting that anger may have a more restricted or individual-specific cortical response pattern compared to other basic emotions.
Figure 7: Group-level statistical T-maps for the anger events. Channels belonging to statistically significant clusters are marked with white circles. Greyed-out topomaps indicate no significant clusters were found for that band.
Fear: Group-level analysis of fear events revealed profound cortical modulations characterized by multi-frequency engagement as shown in Fig. 8. The $\alpha$ band demonstrated the most extensive response, with 57 electrodes (96.6%) forming a single significant cluster ($p=0.001$, $d=-1.105$ to $-0.136$). This widespread $\alpha$ desynchronization spanned nearly the entire cortical surface, reflecting global cortical disinhibition and enhanced neural readiness characteristic of threat detection states. Remarkably, the $\delta$ band showed substantial power increases, with 25 electrodes (42.4%) distributed across two significant clusters ($p=0.010$ and $p=0.026$, $d=0.222$ to $0.638$), primarily in frontal and temporal regions. This extensive delta synchronization, unique among all analyzed emotions, suggests deep engagement of limbic-cortical circuits involved in threat processing and emotional memory consolidation. The $\beta$ band exhibited significant modulation with 20 electrodes (33.9%) forming a single cluster ($p=0.001$, $d=-1.370$ to $0.288$), likely reflecting enhanced cognitive control and motor preparation networks.
In contrast, neither $\theta$ nor $\gamma$ bands showed statistically significant clustered effects after correction for multiple comparisons, distinguishing fear from emotions that engage these higher-frequency oscillations threat response.
Figure 8: Group-level statistical T-maps for the fear events. Channels belonging to statistically significant clusters are marked with white circles. Greyed-out topomaps indicate no significant clusters were found for that band.
Sadness: Group-level analysis of sadness events revealed widespread cortical activation characterized by prominent mid-frequency modulations as shown in Fig. 9. The $\alpha$ band demonstrated extensive desynchronization with 45 out of 59 electrodes (76.3%) forming a single significant cluster ($p=0.004$, $d=-0.772$ to $-0.258$). This widespread $\alpha$ suppression, encompassing frontal, central, temporal, parietal, and occipital regions, indicates comprehensive cortical disinhibition and heightened neural processing associated with sadness. The $\beta$ band showed moderate but significant activation with 12 electrodes (20.3%) forming a cluster ($p=0.005$, $d=-0.623$ to $0.173$), concentrated in parietal and parieto-occipital regions, likely reflecting cognitive-emotional regulation processes. Notably, the $\delta$ band showed highly focal activation with only 1 electrode (1.7%) at F1 reaching significance ($p=0.015$, $d=-0.505$ to $-0.169$), suggesting limited involvement of deep emotional processing circuits compared to high-arousal emotions like fear.
Neither $\theta$ nor $\gamma$ bands showed statistically significant clustered effects after correction for multiple comparisons. The predominant $\alpha$ desynchronization coupled with selective $\beta$ engagement and minimal low-frequency involvement constitutes the characteristic neurophysiological signature of sadness.
Figure 9: Group-level statistical T-maps for the sadness events. Channels belonging to statistically significant clusters are marked with white circles. Greyed-out topomaps indicate no significant clusters were found for that band.
In summary, our research systematically delineates the unique neurophysiological signatures of six basic emotions. A common pattern observed was widespread $\alpha$ band desynchronization, which, acting as a general index of cortical arousal, was present across all emotions. Building upon this, several emotions exhibited characteristic spectral patterns dominated by specific frequency bands: Fear was uniquely defined by a concurrent, profound $\alpha$ band suppression and widespread $\delta$ band synchronization; Happiness was distinguished from other emotions by its hallmark large-scale $\gamma$ band synchronization; while Disgust was identified by its unique marker of localized $\theta$ band suppression in the right posterior region. Meanwhile, other emotions displayed more subtle or mixed patterns: Sadness was characterized by broad desynchronization across both the $\alpha$ and $\beta$ bands; the response elicited by Surprise was more topographically constrained; and Anger produced the weakest and most localized PSD changes, indicating its cortical signature is less distinct under the current paradigm. The ability of our study to consistently resolve these multi-level neural oscillatory patterns—which are also corroborated by existing neuroscience research—ultimately confirms the effectiveness of the FIRMED precise annotation paradigm in capturing and differentiating genuine, fine-grained neurophysiological responses of emotion.
IV-B Autonomic Response Analysis
GSR serves as a sensitive indicator of sympathetic nervous system activity and emotional arousal [47], with SCR being key markers of autonomic activation [48]. We analyzed the occurrence and characteristics of SCRs within the 4-second event windows surrounding the annotated emotional events, comparing these responses to the pre-trial 10-second baseline period. Typically, during the baseline period, the GSR signal often exhibited a gradually declining trend, indicative of a resting or relaxing state.
To illustrate the differential autonomic response patterns across emotional categories, we present representative physiological recordings for emotions with contrasting arousal levels. Fig. 10 displays GSR, PPG, and ECG signals from a fear trial, exemplifying high-arousal emotional responses. In marked contrast to the baseline’s declining GSR trajectory, fear-annotated events consistently triggered distinct SCRs—rapid increases in skin conductance amplitude characteristic of sympathetic activation [49]. Notably, Events 3 and 11 demonstrate pronounced SCRs immediately following the annotated timestamps, while Event 1 exhibits a sustained conductance elevation. Event 7 shows a more moderate SCR, potentially reflecting individual variability in fear intensity perception, yet still deviating significantly from baseline patterns.continuously declining baseline.
Figure 10: The PPG, GSR and ECG signals recorded for a complete trial to the emotion of fear. The upper part of the figure shows the overall change trend of these physiological signals during the experiment: the initial gray background area is the baseline, and the red vertical dotted line is the marked timestamp. The lower part of the figure zooms in on events 1, 3, 7 and 11, showing the details of the physiological signals around the emotional start time recalled by the participant.
In stark contrast, Fig. 11 presents physiological recordings from a sadness trial, representing low-arousal emotional processing. Throughout both baseline and stimulus presentation periods, the GSR signal maintained its characteristic declining trajectory without significant perturbations. Critically, the annotated sadness events failed to elicit discernible SCRs, the GSR signal continued its downward trend uninterrupted. This absence of phasic sympathetic responses during participant-reported emotional moments provides compelling evidence for the arousal-specific nature of autonomic activation patterns.
Figure 11: The PPG, GSR, and ECG signals recorded during a complete experiment for the emotion sadness. The figure shows the overall change trend of these physiological signals during the experiment: the initial gray background area is the baseline, and the red vertical dotted line is the marked timestamp.
To quantify the observed differences in SCR elicitation across all participants and emotion categories, we calculated the GSR increase percentage, the methodology for which is detailed in (2):
$\begin{split}GSR_{increase\%}(e)&=\frac{\sum_{i=1}^{n}\sum_{j=1}^{m_{i}}I(SCR_{ij}(e)>0.05\mu S)}{n\times\overline{m}}\\ &\times 100\%\end{split}$ (2)
[…]
Fig. 12(a) presents the mean values of this SCR-based metric for each of the six basic emotion categorically, ’Fear’ events exhibited the highest GSR increase percentage at 98.8% , followed by ’Surprise’ 92.6%, ’Happiness’ 87.2%, , and ’Anger’ 85.2%, distinctly grouped by high and low arousal levels. As evident from the figure, high-arousal emotions consistently elicited stronger or more frequent SCRs. Conversely, low-arousal emotions ’Sadness’ 7.4%, and ’Disgust’ 4.6% showed substantially lower mean values for this metric, indicating a significantly attenuated phasic sympathetic response. This dichotomy is further emphasized in Fig. 12(b), which directly compares the mean GSR Increase Percentage between the aggregated high-arousal and low-arousal emotion groups. The high-arousal group demonstrated a mean GSR increase percentage of 91.0% , whereas the low-arousal group showed a mean of only 6.0%.
((a)) GSR increase percentage for each of the six basic emotions
((b)) The difference in GSR increase percentage between the high arousal group and the low arousal group
Figure 12: Differential GSR increase percentage to emotions of varying arousal.
These quantitative findings provide robust physiological validation for our annotation paradigm. The temporal alignment between participant-reported emotional events and objective autonomic markers—particularly the reliable SCR elicitation for high-arousal emotions and its absence for low-arousal states—confirms that our immediate recall-based approach successfully captures physiologically meaningful emotional moments.
IV-C Comparative of Immediate versus Delayed Recall Annotation Paradigms
To evaluate the effectiveness of the proposed annotation approach, we conducted a controlled comparative experiment examining the differences between immediate and delayed recall annotation methods. The study employed a within-participant design with the same 20 participants from the FIRMED. We selected 48 videos from the SEED-VII dataset as stimuli, comprising 8 videos for each of the six basic emotions. The experimental protocol was carefully balanced: the first 24 videos utilized the immediate recall annotation paradigm (as in FIRMED), while the remaining 24 videos employed the delayed recall approach (as in SEED-VII). This design allowed us to assess the impact of memory bias on annotation quality through both subjective and objective measures.
First, participants rated their subjective confidence regarding annotation accuracy using a 7-point Likert scale (1 = completely unsure, 7 = extremely certain) [50]. Descriptive statistics indicated that the mean confidence in the immediate annotation condition (Mean = 6.85, SD = 0.21) was numerically higher than in the delayed condition (Mean = 5.91, SD = 1.35). Crucially, there was a significant difference in response variability between the conditions. A Levene’s test [51] was conducted to statistically evaluate the equality of variances, revealing significant heterogeneity (F(1, 38) = 47.2, $p$ <0.001). To further quantify this relative dispersion, we computed the coefficient of variation (CV) [52], which calculated as CV = SD/Mean. The CV for the immediate condition was only 3.1%, substantially lower than the delayed condition’s 22.8%. This suggests that the delayed annotation significantly increased annotation inconsistency and uncertainty.
Second, building on our findings from Section IV-B, which established the SCR as a robust physiological marker of high-arousal emotions, we quantitatively evaluated the temporal concordance between annotated events and autonomic activation. We calculated the probability of valid SCR occurrence within a 4-second window at the annotated event. The results demonstrated that the SCR concordance probability under the immediate annotation paradigm is 76.3% higher than the delayed paradigm 69.7%. This numerical trend aligns with the subjective confidence results, suggesting that the delayed annotation paradigm may impair temporal alignment between labels and actual physiological response peaks. The converging evidence from both subjective confidence ratings and objective physiological concordance robustly supports the efficacy of the immediate replay discrete annotation paradigm in mitigating memory bias and enhancing label accuracy.
The analyses presented in this section provide multifaceted validation for our proposed annotation paradigm. First, through internal physiological validation, we established a strong temporal correspondence between the discrete annotation points within the FIRMED dataset and objective physiological signatures, spanning both emotion-specific neural oscillations and autonomic arousal indexed by SCRs. Furthermore, a comparative experiment demonstrated the superiority of our immediate annotation method over the traditional delayed approach in mitigating key indicators of memory bias, namely subjective uncertainty and response variability. Collectively, these findings confirm that the proposed paradigm offers a foundational solution to the challenge of label noise in emotion research by addressing the issue at the point of data annotation, rather than through subsequent, post-hoc corrections.
V Impact of Annotation on Emotion Recognition
This section presents a computational analysis to determine if our precise annotation paradigm yields superior emotion recognition performance. To isolate the impact of annotation quality, we conducted a rigorously controlled experiment. We created two datasets from the identical underlying physiological recordings: FIRMED, which uses our precise, recall-based annotations, and a control dataset, FIRMED-Whole, which applies traditional whole-trial labeling. By keeping all other experimental factors—such as feature extraction, classification models, and cross-validation procedures—identical across both conditions, this comparative framework allows any resulting performance differences to be robustly attributed to the annotation strategy itself.
V-A Dataset Construction
In order to achieve comparative experiments, this study constructed two datasets as follows:
FIRMED: For this dataset, we implemented a conservative data augmentation strategy that follows the temporal precision of our annotations while acknowledging the inherent uncertainty in emotion onset timing. For each annotated emotional event at time $t_{event}$, we extracted three overlapping 4-second windows: the primary window centered at $t_{event}$ ([ $t_{event}$ 2s, $t_{event}$ + 2s]), and two adjacent windows shifted by ±1 second ([ $t_{event}$ 3s, $t_{event}$ + 1s] and [ $t_{event}$ 1s, $t_{event}$ + 3s]). This approach ensures that each window maintains substantial overlap with the annotated event time, preserving the temporal precision that distinguishes our annotation paradigm while providing modest data augmentation. All windows derived from the same emotional event inherited the original emotion label.
FIRMED-Whole: To simulate traditional whole-trial annotation, we used the entire physiological recording during the presentation of each video stimulus as described in Section III-C. This entire about 1-minute segment was assigned the single emotion label reported by the participant at the timestamp within that video. For the FIRMED-Whole dataset, to maintain consistency in segment length with FIRMED for feature extraction and model input, and also to increase the sample size, we also performed overlapping 4-second window segmentation on whole data, with a step of 2-second.
V-B Feature Extraction
For all generated 4-second data segments in both the FIRMED and FIRMED-Whole datasets, we extracted the following features from the preprocessed data:
For EEG data, we selected Differential Entropy (DE) [53], a feature well-established for its efficacy in capturing the complex, non-linear dynamics of neural oscillations in emotion recognition. Specifically, DE was computed for each of the 59 EEG channels across five standard frequency bands, resulting in a 295-dimensional feature vector (59 channels × 5 bands) for each 4-second segment. The calculation method is shown in equation (3):
$DE=\frac{1}{2}\log(2\pi e\sigma^{2})$ (3)
where $\sigma^{2}$ is the variance of the signal in each frequency band, obtained via bandpass filtering.
From the GSR data, we extracted the skewness of its first derivative (4), a choice motivated by its ability to precisely quantify the rapid, phasic increases in skin conductance that our physiological analysis identified as a hallmark of high-arousal emotions [54].
$Skewness=\frac{\frac{1}{N}\sum_{i=1}^{N}\left(v_{i}-\mu_{v}\right)^{3}}{\sigma_{v}^{3}}$ (4)
where $v_{i}$ denotes the GSR derivative at time point $i$, representing the instantaneous rate of change in skin conductance. $N$ is the total number of samples within the 4-second analysis window. $\mu_{v}$ represents the mean of all derivative values across the window, indicating the average trend of GSR change. $\sigma_{v}$ is the standard deviation of the derivative values, measuring the variability of GSR change rates throughout the window.
From ECG data, we computed Root Mean Square of Successive Differences (RMSSD) [55] to quantify short-term heart rate variability:
$RMSSD=\sqrt{\frac{1}{N-1}\sum_{i=1}^{N-1}(RR_{i+1}-RR_{i})^{2}}$ (5)
where $RR_{i}$ represents the $i$-th R-R interval in seconds, and $N$ is the total number of R-R intervals within the 4-second window. This metric is particularly sensitive to parasympathetic nervous system activity, with lower values indicating sympathetic dominance during high-arousal states and higher values reflecting parasympathetic influence in calm states.
From PPG data, we calculated the change rate of Pulse Wave Amplitude (PWA) [56] to capture peripheral vascular dynamics:
${PWA}_{i}={PPG}_{{peak},i}-{PPG}_{{valley},i}$ (6)
$\Delta PWA=\frac{\overline{PWA}_{post}-\overline{PWA}_{pre}}{\overline{PWA}_{total}}$ (7)
where ${PPG}_{{peak},i}$ and $PPG_{valley,i}$ represent the systolic peak and diastolic valley of the $i$-th pulse wave, respectively. $\overline{PWA}_{{pre}}$ and $\overline{PWA}_{post}$ denote the mean amplitudes in the first and second halves of the window, while $\overline{PWA}_{total}$ is the overall mean. This feature detects emotion-induced vasoconstriction or vasodilation.
V-C Classification Results and Analysis
V-C1 EEG-based Emotion Recognition
To comprehensively evaluate the impact of annotation precision on emotion recognition performance, we employed multiple classification algorithms spanning machine learning and deep learning approaches. We selected eight EEG-based representative classifiers to ensure our findings are robust across different algorithmic, as follows:
Support Vector Machine (SVM) [57]: RBF kernel with hyperparameters optimized via grid search ($C\in{0.1,1,10,100}$, $\gamma\in{0.001,0.01,0.1,1}$)
k-Nearest Neighbors (k-NN) [58]: k=5, using Euclidean distance metric.
Multi-Layer Perceptron (MLP) [59]: Three hidden layers (128, 64, 32 neurons) with ReLU activation and dropout (0.3).
EEGNet [60]: A compact Convolutional Neural Network (CNN) architecture specifically designed for EEG signals, effectively employing depthwise separable convolutions to learn spatio-temporal features..
Long Short-Term Memory (LSTM) [61]: A two-layer stacked LSTM architecture was used to capture temporal dependencies within the EEG sequences.
Dynamic Graph Convolutional Neural Network (DGCNN) [62]: A graph neural network variant where the graph’s adjacency matrix is treated as learnable parameters, allowing for dynamic capture of inter-channel relationships.
Multi-domain Dynamic Graph Representation Learning (MD2GRL) [63]: A method designed for learning dynamic graph representations from EEG by simultaneously leveraging features from both time and frequency domains.
Transformer [64]: Self-attention based model adapted for EEG sequences with positional encoding and 4 attention heads.
To assess cross-subject generalization, we used a Leave-One-Subject-Out (LOSO) cross-validation scheme. For each of the 20 participants, the model was trained on the other 19 and tested on the held-out individual. Final performance metrics were then averaged across all 20 test results.
Table II presents the comprehensive performance comparison between FIRMED and FIRMED-Whole datasets. It can be seen that all algorithms have improved in classification accuracy, with an average improvement of about 9.7%, among which the best model DGCNN has an improvement of 6.2%, the confusion matrix of DGCNN in the two datasets is shown in the Fig. 13. The average improvement of above algorithms in different emotion classifications is shown in Table III.
TABLE II: Performance of eight classification algorithms on EEG data with two annotation methods
Classifier Dataset Accuracy Precision Recall F1-score
SVM FIRMED $45.2\pm 4.8$ $43.6\pm 5.1$ $44.8\pm 4.9$ $44.2\pm 5.0$
FIRMED-Whole $31.1\pm 5.2$ $30.5\pm 5.5$ $30.7\pm 5.3$ $30.1\pm 5.4$
k-NN FIRMED $49.7\pm 4.3$ $47.2\pm 4.6$ $48.3\pm 4.4$ $47.7\pm 4.5$
FIRMED-Whole $40.5\pm 5.0$ $38.0\pm 5.3$ $39.1\pm 5.1$ $39.5\pm 5.2$
MLP FIRMED $56.8\pm 3.9$ $55.4\pm 4.2$ $56.4\pm 4.0$ $55.9\pm 4.1$
FIRMED-Whole $48.6\pm 4.5$ $49.1\pm 4.8$ $50.2\pm 4.6$ $49.6\pm 4.7$
EEGNet FIRMED $57.6\pm 3.5$ $56.3\pm 3.8$ $55.2\pm 3.6$ $55.7\pm 3.7$
FIRMED-Whole $48.4\pm 4.1$ $47.0\pm 4.4$ $47.0\pm 4.2$ $46.5\pm 4.3$
LSTM FIRMED $54.1\pm 3.7$ $52.8\pm 4.0$ $53.7\pm 3.8$ $53.2\pm 3.9$
FIRMED-Whole $41.9\pm 4.3$ $40.5\pm 4.6$ $41.5\pm 4.4$ $41.0\pm 4.5$
MD2GRL FIRMED $58.9\pm 3.2$ $57.6\pm 3.5$ $58.5\pm 3.3$ $58.0\pm 3.4$
FIRMED-Whole $52.7\pm 3.8$ $51.3\pm 4.1$ $50.3\pm 3.9$ $50.8\pm 4.0$
DGCNN FIRMED $\mathbf{61.4\pm 3.1}$ $\mathbf{60.1\pm 3.4}$ $\mathbf{61.0\pm 3.2}$ $\mathbf{60.5\pm 3.3}$
FIRMED-Whole $55.2\pm 3.6$ $53.8\pm 3.9$ $52.9\pm 3.7$ $52.3\pm 3.8$
Transformer FIRMED $54.8\pm 3.6$ $53.5\pm 3.9$ $54.4\pm 3.7$ $53.9\pm 3.8$
FIRMED-Whole $42.6\pm 4.2$ $41.2\pm 4.5$ $42.2\pm 4.3$ $41.7\pm 4.4$
Figure 13: The confusion matrix of DGCNN in the two datasets, where the left one is FIRMED and the right one is FIRMED-Whole.
TABLE III: The average accuracy and improvement of each emotion in different classifiers (%)
Emotion FIRMED FIRMED-Whole Improvement
Fear $62.8$ $52.5$ $+10.3$
Happiness $60.1$ $51.0$ $+9.1$
Surprise $56.6$ $46.8$ $+9.8$
Sadness $53.4$ $40.2$ $+13.2$
Disgust $49.9$ $41.6$ $+8.3$
Anger $46.2$ $38.6$ $+7.6$
Average $\mathbf{54.8}$ $\mathbf{45.1}$ $+\mathbf{9.7}$
The analysis of these results yields several key insights. First, the superior performance of graph-based models DGCNN and MD2GRL validates the importance of capturing inter-channel relationships in EEG data. Second, universal performance gains were observed across all algorithms, highlighting the algorithm-agnostic benefit of high-precision annotation. Furthermore, the performance delta between traditional and deep learning approaches widened on the FIRMED dataset, indicating that high-quality labels are crucial for complex models to learn robust and meaningful patterns. Finally, the highest classification accuracies were achieved for ”Fear” and ”Happiness,” a finding that directly corroborates our conclusion from the physiological analysis that these two emotions have the most unique neural signal signatures.
V-C2 Multimodal Emotion Recognition
Given that the DGCNN demonstrated superior performance among the evaluated models in the EEG-based emotion recognition tasks, it was selected as the primary architecture for a detailed exploration of multimodal fusion. To effectively integrate features derived from peripheral physiological signals (GSR, ECG, PPG) with the graph-based representations learned from EEG data by the DGCNN, we adopted a feature-level fusion strategy. This approach allows each data modality to be processed by an architecture best suited to its inherent structure before their learned representations are combined for final classification.
The multimodal DGCNN architecture was conceptualized with two primary processing streams, the outputs of which were combined differently depending on the specific modality combination under evaluation, as presented in Table IV.
The EEG data, these graph convolutional layers from DGCNN are designed to extract high-level feature embeddings that capture both channel-specific activities and complex spatio-temporal dependencies. The output from this stream, subsequent to the graph convolution and any graph pooling operations, is a flattened feature vector $F_{EEG}$ representing the learned EEG-specific information. For each specific multimodal combination to be evaluated, a feature vector was constructed using only the features from the peripheral signal(s) pertinent to that combination (Skewness, RMSSD, $\Delta$PWA, as defined in Section V-B). This tailored peripheral feature vector was then processed through a MLP sub-network, consisting of one dense layer with a non-linear activation function, to yield an abstract peripheral feature representation, denoted as $F_{Peripheral}$.
In all multimodal scenarios, the learned EEG feature vector $F_{EEG}$ was concatenated with the corresponding abstract peripheral feature vector $F_{Peripheral}$, which represented the specific peripheral signal(s) chosen for that particular test combination. This comprehensive vector was then fed into identical final fully-connected layers, culminating in a softmax layer, for predicting the target emotional states.
TABLE IV: Performance Comparison of Different EEG-based Modality Combinations using DGCNN (%)
Modality Combination Dataset Accuracy Precision Recall F1-score
EEG-only FIRMED $61.4\pm 3.1$ $60.1\pm 3.4$ $61.0\pm 3.2$ $60.5\pm 3.3$
FIRMED-Whole $55.2\pm 3.6$ $53.8\pm 3.9$ $52.9\pm 3.7$ $52.3\pm 3.8$
EEG + PPG FIRMED $63.8\pm 3.2$ $62.6\pm 3.5$ $63.2\pm 3.3$ $62.9\pm 3.4$
FIRMED-Whole $56.2\pm 3.7$ $56.0\pm 4.0$ $56.6\pm 3.8$ $56.3\pm 3.9$
EEG + ECG FIRMED $64.5\pm 3.1$ $63.3\pm 3.4$ $64.0\pm 3.2$ $63.6\pm 3.3$
FIRMED-Whole $58.0\pm 3.6$ $56.8\pm 3.9$ $57.5\pm 3.7$ $57.1\pm 3.8$
EEG + GSR FIRMED $67.0\pm 2.9$ $65.8\pm 3.1$ $66.5\pm 3.0$ $66.1\pm 3.0$
FIRMED-Whole $60.0\pm 3.4$ $58.8\pm 3.6$ $59.5\pm 3.5$ $59.1\pm 3.5$
EEG + GSR + ECG + PPG FIRMED $\mathbf{69.5\pm 2.8}$ $\mathbf{68.3\pm 3.0}$ $\mathbf{69.0\pm 2.9}$ $\mathbf{68.6\pm 2.9}$
FIRMED-Whole $\mathbf{63.5\pm 3.3}$ $\mathbf{61.3\pm 3.5}$ $\mathbf{62.0\pm 3.4}$ $\mathbf{61.6\pm 3.4}$
The results for the DGCNN model, presented in Table IV, unequivocally demonstrate the benefits of integrating peripheral physiological signals with EEG data for emotion recognition. All tested multimodal combinations surpassed the EEG-only baseline performance on both FIRMED and FIRMED-Whole datasets. Notably, the comprehensive fusion of EEG with all peripheral signals yielded the highest accuracy, achieving 69.5% on the FIRMED dataset. Among the single peripheral additions, EEG+GSR provided the most substantial improvement over EEG-only, reaching an accuracy of 67.0%, underscoring GSR’s strong capability in capturing arousal-related emotional dynamic.
Critically, the advantage of precise temporal annotation remained consistent and significant across all multimodal configurations. For every combination, models trained on the FIRMED dataset achieved superior accuracy compared to those trained on FIRMED-Whole; for example, the full multimodal accuracy on FIRMED was 6.0% higher than its counterpart on FIRMED-Whole. This robustly illustrates that the enhanced data quality derived from our precise annotation paradigm effectively synergizes with the richer information from combined physiological signals, leading to optimal emotion recognition performance and further validating the overall dataset construction and annotation methodology.
VI Conclusion
To address the prevalent challenges of label noise in video-induced emotion datasets, this paper introduces FIRMED, a multimodal dataset developed using an immediate recall-based annotation paradigm. Our comprehensive validation demonstrates the paradigm’s methodological rigor and empirical effectiveness across multiple analytical dimensions.
Physiologically, we established temporal alignment between participant-reported emotional events and objective neurophysiological signatures, including emotion-specific oscillatory patterns across EEG frequency bands and distinctive autonomic activation profiles. This provides validation of the subjective-objective correspondence that has long been theoretically postulated but rarely systematically verified in emotion research. Computationally, classification models trained on precisely annotated data demonstrated consistent and substantial performance improvements across eight distinct algorithms, and multimodal approaches demonstrated superior performance compared to EEG-only configurations. These computational findings align with conclusions from SEED-VII [21], reinforcing that annotation quality serves as a critical determinant of recognition performance.
We acknowledge several limitations that warrant consideration. The dataset scale constrains statistical power and generalizability across diverse populations. The fixed 4-second analysis window, while physiologically informed, may not optimally capture the heterogeneous temporal dynamics of emotional responses across individuals and emotional contexts. Future investigations should prioritize expanding participant diversity across demographics, implementing adaptive temporal windowing techniques that accommodate individual response variability.

### v2

@@
 Abstract
-Traditional video-induced emotion physiological datasets often use whole-trial annotation, assigning a single emotion label to all data collected during an entire trial.
-This coarse-grained annotation approach misaligns with the dynamic and temporally localized nature of emotional responses as they unfold with video narratives, introducing label noise that limits emotion recognition algorithm evaluation and performance.
-To solve the label noise problem caused by coarse-grained annotation, we propose a fine-grained annotation method through an immediate recall paradigm.
-This paradigm integrates an immediate video replay phase after the initial stimulus viewing, allowing participants to precisely mark the onset timestamp, emotion label, and intensity based on their immediate recall.
+Traditional video-induced emotion physiological datasets typically employ whole-trial annotation, assigning a single emotion label to all data collected throughout the experiment.
+This coarse-grained annotation method is inconsistent with the dynamic and temporally localized nature of emotional responses within video narratives, introducing label noise.
+To address the label noise problem caused by coarse-grained annotation, inspired by the Peak-End Rule, we propose a fine-grained annotation method based on an immediate recall paradigm.
+This paradigm incorporates a video replay phase immediately after the initial stimulus viewing, allowing participants to precisely annotate the timestamp, emotion label, and intensity of emotional outbursts based on immediate recall.
 We validate this paradigm through physiological evidence and recognition performance.
-Physiological validation of multimodal signals within participant-marked windows revealed rhythm-specific EEG patterns and arousal-dependent GSR responses—with SCRs appearing in 91% of high-arousal versus 6% of low-arousal emotion windows.
+Physiological validation of multimodal signals within participant-marked windows revealed rhythm-specific EEG patterns; for instance, fear exhibits the most widespread $\alpha$ desynchronization, while happiness shows significant $\gamma$ and $\beta$ enhancement.
+Meanwhile, arousal-dependent GSR responses were observed—with SCRs appearing in 79.8% of high-arousal versus 12.7% of low-arousal emotion windows.
 These objective physiological data changes strongly aligned with subjective annotations, confirming annotation precision.
-For recognition performance, classification experiments showed that models trained on fine-grained annotations achieved 9.7% higher accuracy than traditional whole-trial labeling, despite using less data.
+For recognition performance, classification experiments showed that models trained on fine-grained annotations achieved 9.1% higher accuracy than traditional whole-trial labeling, despite using less data.
 This work not only addresses label noise through fine-grained annotation but also demonstrates that annotation precision outweighs data scale in determining emotion recognition performance.
-Index Terms:
-Emotion Annotation, Label Noise, Multimodal Emotion Dataset, Physiological Validation, Emotion Recognition.
 I Introduction
-Emotion is a subjective experience elicited by external stimuli [1], typically accompanied by physiological responses and behavioral tendencies.
-Emotions play a crucial role in individual survival, social interaction [2], and decision-making processes [3].
+Emotion is a subjective experience elicited by external stimuli [frijda2005emotion], typically accompanied by physiological responses and behavioral tendencies.
+Emotions play a crucial role in individual survival[rolls2014emotion].
 Two primary theoretical frameworks describe emotions: discrete and dimensional models.
-The discrete model, exemplified by Ekman’s seminal work [4], identifies six basic emotions—happiness, sadness, fear, anger, surprise, and disgust—each linked to specific neural circuits.
+The discrete model, exemplified by Ekman’s seminal work [ekman1999basic], identifies six basic emotions—happiness, sadness, fear, anger, surprise, and disgust—each linked to specific neural circuits.
 For instance, fear demonstrates strong associations with amygdala activation.
-In contrast, the dimensional model, represented by Russell’s Circumplex Model [5], characterizes emotions along two dimensions: valence and arousal.
+In contrast, the dimensional model, represented by Russell’s Circumplex Model [russell1980circumplex], characterizes emotions along two dimensions: valence and arousal.
 These complementary frameworks establish a robust theoretical foundation for emotion research.
 To explore neural mechanisms and develop recognition methods for emotions, researchers have designed various experimental paradigms to simulate and capture human emotional responses.
-These experiments typically employ music [6], images [7], or videos [8] as stimuli while collecting participants’ physiological and behavioral signals.
-Among these modalities, video stimuli demonstrate superior performance in eliciting complex emotions due to their dynamic nature and multisensory input, thereby offering higher ecological validity compared to static images or single-modality audio stimuli [9].
-Data collection encompasses both physiological signals (e.g., electroencephalogram (EEG), electrocardiogram (ECG)) [10, 11] and non-physiological signals (e.g., facial expressions) [12], providing comprehensive documentation of emotional manifestations.
-Physiological signals offer high temporal resolution and objectivity, enabling precise capture of neural and physiological changes associated with emotions [13].
-Non-physiological signals provide behavioral cues that complement physiological measures, further deepening our multidimensional understanding of emotional mechanisms [14].
+These experiments typically employ music [eerola2012review], images [junghofer2001fleeting], or videos [soleymani2011multimodal] as stimuli while collecting participants’ physiological and behavioral signals.
+Among these modalities, video stimuli demonstrate superior performance in eliciting complex emotions due to their dynamic nature and multisensory input, thereby offering higher ecological validity compared to static images or single-modality audio stimuli [siedlecka2019experimental].
+Data collection encompasses both physiological signals (e.g., electroencephalogram (EEG), electrocardiogram (ECG)) [alarcao2017emotions, hsu2017automatic, shen2025physiological] and non-physiological signals (e.g., facial expressions) [liu2017facial], providing comprehensive documentation of emotional manifestations.
+Physiological signals offer high temporal resolution and objectivity, enabling precise capture of neural and physiological changes associated with emotions [shu2018review].
+Non-physiological signals provide behavioral cues that complement physiological measures, further deepening our multidimensional understanding of emotional mechanisms [pillalamarri2025review].
 Despite continuous developments in experimental paradigms for emotion research, existing methods face significant challenges in emotion annotation accuracy.
-Datasets such as SEED [15] and DEAP [16] employ coarse-grained annotation strategies, assigning single emotion labels to entire multi-minute trials.
+Datasets such as SEED [zheng2015investigating] and DEAP [koelstra2011deap] employ coarse-grained annotation strategies, assigning single emotion labels to entire multi-minute trials.
 While this approach reduces annotation time, it overlooks the temporal instability and discrete nature of emotions, resulting in substantial label noise.
 To address this limitation, Zhang et al.
-[17] proposed an emotion localization task for the SEED dataset.
+[zhang2022eeg] proposed an emotion localization task for the SEED dataset.
 They segmented each video stimulus into multiple intervals and recruited 20 annotators to evaluate whether each segment triggered the target emotion.
 Segments receiving majority agreement were labeled as ground truth, thereby reducing noise by focusing on highly relevant emotional segments.
-Building upon this foundation [17], Zhang et al.
-[18] introduced TAS-Net, an unsupervised time-aware sampling network based on deep reinforcement learning.
+Building upon this foundation [zhang2022eeg], Zhang et al.
+[zhang2023unsupervised] introduced TAS-Net, an unsupervised time-aware sampling network based on deep reinforcement learning.
 This approach automatically detects key emotional segments in EEG signals while reducing labeling noise from traditional fixed-label methods.
@@
 Similarly, Jiang et al.
-proposed REMONET [19], which mitigates sentiment label noise through a multi-regularization co-learning framework that coordinates different self-supervised data views to prevent model overfitting to noisy labels.
-However, these post-hoc approaches—whether through manual re-annotation or deep learning techniques—do not fundamentally resolve the inherent ambiguity in mapping physiological data to subjective emotion labels [20].
-In recent efforts to address annotation granularity, researchers have developed multimodal datasets such as SEED-VII [21] and MGEED [22], aiming to achieve finer-grained emotional labeling.
-SEED-VII represents a significant advancement by incorporating both trial-by-trial intensity ratings and post-session continuous annotation.
-Specifically, participants provide immediate intensity feedback after each trial, then retrospectively assign frame-level emotional intensities while reviewing all videos at the session’s conclusion.
-This dual-annotation approach attempts to capture both immediate impressions and detailed temporal dynamics.
-Similarly, MGEED employs a comprehensive frame-by-frame retrospective protocol, where participants meticulously annotate their emotional states for each video frame after viewing all experimental stimuli.
-While these methodologies achieve fine-grained temporal annotation, both datasets fundamentally rely on delayed recall mechanisms.
-This “watch first, annotate later” paradigm, despite its temporal granularity, introduces significant cognitive load [23], as participants must accurately remember and differentiate emotional experiences across multiple stimuli viewed over extended periods.
-Moreover, inherent memory bias [24] in delayed reporting may lead to temporal displacement of emotional events or conflation of distinct emotional episodes, potentially compromising annotation reliability.
-To achieve fine-grained emotion annotation while mitigating memory bias, we propose an immediate recall annotation paradigm and establish the Fine-grained Immediate Recall-based Multimodal Emotion Dataset (FIRMED).
-This approach introduces three key methodological advancements:
-1) Precise Temporal Localization via Immediate Recall: To achieve precise temporal localization of emotional events within each trial, our paradigm incorporates an immediate replay phase following stimulus presentation.
-During post-stimulus replay, participants retrospectively identify and mark the discrete time point $t_{\text{event}}$ at which they recall first experiencing the target emotion during initial viewing.
-A fixed-duration window (4 seconds) centered at $t_{\text{event}}$ is then extracted for analysis, associating data within this window with the reported emotion label and intensity.
-This recall-based discrete annotation during immediate replay aims to overcome temporal ambiguity in whole-trial labeling while reducing memory decay biases associated with delayed annotation protocols.
-2) Expanded Multimodal Data for Enhanced Reliability: To improve annotation reliability and objectivity, we expanded the range of collected physiological signals to include EEG, ECG, galvanic skin response (GSR), and photoplethysmogram (PPG) data.
-Additionally, we captured participants’ facial expressions via camera during stimulus viewing, providing behavioral data on emotional responses.
-This multimodal collection strategy enriches data variety and enables cross-validation of subjective annotations using various physiological data.
-3) Ecologically Valid Stimuli: We extensively utilize short videos from social media platforms as stimuli, encompassing authentic scenes including daily life recordings and news reports.
-Compared to traditional movie clips or music videos, these stimuli more accurately reflect natural emotional experiences, offering enhanced ecological validity.
-Participant feedback regarding stimulus effectiveness further supports this choice.
+proposed REMONET [jiang2024remonet], which mitigates sentiment label noise through a multi-regularization co-learning framework that coordinates different self-supervised data views to prevent model overfitting to noisy labels.
+However, these post-hoc approaches—whether through manual re-annotation or deep learning techniques—do not fundamentally resolve the inherent ambiguity in mapping physiological data to subjective emotion labels [winter1997individual].
+Recent efforts like SEED-VII [jiang2024seed] and MGEED [wang2023mgeed] improved temporal granularity using delayed, continuous, or frame-by-frame annotations.
+However, these “watch first, annotate later” paradigms fundamentally rely on delayed recall over extended periods, incurring vulnerability to memory bias [thomas1990memory, colombo2020affect].
+This limitation stems from the human mechanism of emotional memory.
+Research by Daniel Kahneman, particularly the Peak-End Rule, posits that retrospective memory disproportionately retains ’snapshots’ of peak and end moments, rather than a continuous, faithful recording [kahneman1993more, fredrickson1993duration, horwitz2024review].
+Consequently, demanding continuous retrospective annotation forces reconstruction from an inherently biased memory system, potentially more insidious form of label noise despite the fine granularity.
+To overcome the twin challenges of temporal imprecision in coarse-grained methods and memory bias in delayed fine-grained methods, this paper proposes a novel immediate recall annotation paradigm specifically designed to capture high-fidelity emotional moments.
+We introduce the Fine-grained Immediate Recall-based Multimodal Emotion Dataset (FIRMED), created using this paradigm.
+By incorporating an immediate replay phase after each stimulus, our method empowers participants to precisely identify and mark the discrete timestamp ($t_{\text{event}}$) of subjectively experienced emotional peaks based on fresh recall.
+This directly targets the most robustly encoded ’snapshots’ suggested by the Peak-End Rule, aiming to minimize label noise at the source.
+We further bolster the quality and utility of FIRMED through:
+1) Expanded Multimodal Recordings: Synchronized EEG, ECG, Galvanic Skin Response (GSR), photoplethysmogram (PPG), and facial video enable robust cross-validation between subjective labels and objective physiological responses.
+2) Ecologically Valid Stimuli: Traditional datasets using film clips may introduce a cognitive load of narrative comprehension [kimura2019emotional, guo2022emotional].
+In contrast, we selected short videos from social media as our stimuli.
+Most of these videos, consisting of news reports or daily life recordings, can be understood without imposing significant extra cognitive load.
 To validate this paradigm’s effectiveness, we conducted systematic physiological and computational analyses.
-Physiologically, by analyzing multimodal signals within precisely annotated windows, we found strong evidence of both central [25] and autonomic nervous system engagement [26].
-This included discovering emotion-specific spectral patterns in EEG signals and strong temporal coupling between discrete annotations and objective autonomic nervous system activation markers, exemplified by specific skin conductance response (SCR) patterns reliably occurring for high-arousal emotions [27].
+Physiologically, by analyzing multimodal signals within precisely annotated windows, we found strong evidence of both central [hagemann2003central] and autonomic nervous system engagement [kreibig2010autonomic].
+This included discovering emotion-specific spectral patterns in EEG signals and strong temporal coupling between discrete annotations and objective autonomic nervous system activation markers, exemplified by specific skin conductance response (SCR) patterns reliably occurring for high-arousal emotions [xia2017salience].
 Computationally, emotion classification experiments compared the effectiveness of the two labeling strategies.
@@
 II Related Work
-II-A Evolution of Physiological Signal-Based Emotion Datasets
-Video-based emotion elicitation has established itself as the predominant stimulus modality for inducing emotional responses in controlled experimental environments, driving the proliferation of multimodal datasets across the affective computing domain [28].
-These datasets systematically collect multiple physiological signals to provide comprehensive documentation of participants’ emotional states [29].
-Early and highly influential datasets, including DEAP [16] and the foundational SEED series [15, 30, 31], established the methodological foundation for physiological emotion analysis through rich signal acquisition protocols.
+II-A Evolution of video-induced Physiological Signal-Based Emotion Datasets
+Video-induced emotion elicitation has established itself as the predominant stimulus modality for inducing emotional responses in controlled experimental environments [pei2025identifying], driving the proliferation of multimodal datasets across the affective computing domain [li2022eeg].
+These datasets systematically collect multiple physiological signals to provide comprehensive documentation of participants’ emotional states [wang2023unlocking].
+Early and highly influential datasets, including DEAP [koelstra2011deap] and the foundational SEED series [zheng2015investigating, zheng2018emotionmeter, liu2021comparing], established the methodological foundation for physiological emotion analysis through rich signal acquisition protocols.
 These seminal works adopted a standardized experimental paradigm involving presentation of emotionally evocative stimuli—typically music videos or film clips—followed by collection of self-reported emotional assessments.
@@
 This coarse-grained annotation strategy ignored the inherently dynamic nature of video stimuli.
-Emotional responses are characteristically triggered by specific narrative elements—including pivotal plot developments, climactic sequences, or sudden scene transitions—rather than being uniformly sustained throughout extended video segments [32].
-Consequently, this labeling granularity introduces label noise [19], as temporal segments that fail to elicit the target emotion become erroneously associated with that emotional state.
+Emotional responses are characteristically triggered by specific narrative elements—including pivotal plot developments, climactic sequences, or sudden scene transitions—rather than being uniformly sustained throughout extended video segments [hakim2013computational].
+Consequently, this labeling granularity introduces label noise [jiang2024remonet], as temporal segments that fail to elicit the target emotion become erroneously associated with that emotional state.
 In direct response to these temporal granularity limitations, recent dataset development efforts have pursued increasingly sophisticated annotation methodologies.
-The SEED-VII dataset [21] introduced a delayed continuous rating paradigm representing a significant methodological advancement.
-In this approach, participants retrospectively review all trial videos upon session completion and provide continuous intensity annotations.
-Similarly, the MGEED database [22] implemented a delayed frame-by-frame self-assessment protocol, requiring participants to annotate their emotional states retrospectively following complete viewing session.
+Datasets such as SEED-VII [jiang2024seed] and MGEED [wang2023mgeed] represent significant advancements, employing delayed recall paradigms like continuous ratings or frame-by-frame assessments performed after the entire experimental session to achieve fine-grained labeling.
+While these approaches improve temporal resolution compared to whole-trial methods, they introduce a critical challenge: memory bias.
+Research of the Peak-End Rule, suggests that our retrospective memory of an affective experience is not a continuous, faithful recording but is disproportionately dominated by its most intense moment (the “peak”) and its final moment (the “end”) [kahneman1993more, fredrickson1993duration].
+Therefore, asking participants to retrospectively draw a continuous intensity curve forces them to reconstruct a timeline their memory has not accurately stored, potentially introducing label noise despite the fine granularity.
 To systematically characterize these dataset developments, Table I provides a comprehensive comparison of representative emotion datasets.
 The analysis reveals a clear evolutionary trajectory from coarse whole-trial labeling toward increasingly fine-grained temporal annotations.
+It can be seen that exploring more effective emotion labeling paradigms is still a key issue worthy of research in affective computing.
 TABLE I: Evolution of Multimodal Emotion Datasets: Labeling Granularity Paradigm Progression
@@
 Coarse-grained annotation paradigm
-2011 MAHNOB-HCI [8] N=27
+2011 MAHNOB-HCI [soleymani2011multimodal] N=27
 20 video clips
@@
 (Single label per trial)
-2012 DEAP [16] N=32
+2012 DEAP [koelstra2011deap] N=32
 40 music videos (60s)
@@
 (Single label per trial)
-2013 SEED [15] N=15
+2013 SEED [zheng2015investigating] N=15
 45 clips ($\sim$4 min)
@@
 (Single label per trial)
-2017 DREAMER [33] N=23
+2017 DREAMER [katsigiannis2017dreamer] N=23
 18 clips (60s)
@@
 (Single label per trial)
-2018 AMIGOS [34] N=40
+2018 AMIGOS [miranda2018amigos] N=40
 20 videos
@@
 (Single label per trial)
-2019 MPED [35] N=23
+2019 MPED [song2019mped] N=23
 28 clips (2.5-5 min)
@@
 (Single label per trial)
-2021 SEED-V [31] N=20
+2021 SEED-V [liu2021comparing] N=20
 45 clips (2-4 min)
@@
 Fine-grained annotation paradigm
-2023 MGEED [22] N=17
+2023 MGEED [wang2023mgeed] N=17
 9 clips ($\sim$4 min)
@@
 (Delayed post-session)
-2024 SEED-VII [21] N=20
+2024 SEED-VII [jiang2024seed] N=20
 80 clips (2-5 min)
 4 sessions
-EEG (62-ch), EOG 6 Discrete emotions + Neutral Continuous
+EEG (62-ch), EOG 6 Discrete emotions + Neutral Emotion category and intensity labels for each 4-s time windows
 (Delayed post-session)
-Abbreviations: VAD = Valence-Arousal-Dominance; RSP = Respiration; BVP = Blood Volume Pulse; Temp. = Temperature; A-V = Audio-Visual
-II-B Challenges of Emotion Annotation
-The systematic adoption of whole-trial labeling across early multimodal emotion datasets [16, 15, 34, 33, 8, 35].
-This annotation approach suffers from temporal imprecision: emotions constitute dynamic phenomena that evolve synchronously with video narratives [32], exhibiting rapid fluctuations and transitions rather than maintaining static states across extended durations.
-The assignment of uniform labels to multi-minute segments obscures emotional transitions and fluctuations [36], causing label noise to be structurally embedded in the dataset.
-While recent datasets such as SEED-VII and MGEED have achieved significant improvements in labeling granularity through delayed recall paradigms.
-These approaches require participants to complete entire experimental sessions—encompassing numerous stimuli viewed over extended periods (typically 30 minutes to over 1 hour)—before retrospectively annotating their emotional states across all preceding trials.
-This delayed annotation strategy, despite achieving fine temporal granularity, it also increases the risk of memory bias, which may reduce the reliability of annotations.
-Cognitive psychology research [37] demonstrates that episodic memory for specific emotional events and their nuanced temporal characteristics becomes increasingly unreliable when recall is required after extended delays involving multiple distinct experiences.
-The persistence of these limitations—temporal imprecision in whole-trial labeling and memory bias in delayed fine-grained annotation—underscores the urgent need for innovative annotation paradigms capable of achieving fine temporal granularity while maintaining annotation reliability.
+Legend: VAD = Valence-Arousal-Dominance; RSP = Respiration; BVP = Blood Volume Pulse; Temp. = Temperature; A-V = Audio-Visual
+II-B Emotion Recognition Based on Physiological Signals
+Alongside the evolution of datasets, the algorithms for classifying emotions from physiological signals have also advanced significantly.
+This development has progressed along two primary tracks: single-modality analysis, which focuses on rich signals like EEG, and multimodal fusion, which integrates information from multiple physiological sources to achieve a more comprehensive understanding of the emotional state[chen2023multi].
+In the single-modality (primarily EEG-based) track, the field has transitioned from traditional machine learning methods like Support Vector Machines (SVM) [wang2011eeg] and k-Nearest Neighbors (k-NN) [li2018emotion]—which often rely on hand-crafted features like Differential Entropy (DE) [duan2013differential]—to end-to-end deep learning.
+Architectures such as EEGNet [lawhern2018eegnet] capture spatio-temporal features, while Recurrent Neural Networks (RNNs) like Long Short-Term Memory (LSTM) [graves2012long] model the temporal dependencies.
+More recently, Graph Convolutional Neural Networks (GCNs) have shown strong performance by modeling the brain’s functional connectivity as a dynamic graph [song2018eeg, tang2024multi].
+However, emotions are complex, multi-system phenomena, involving both the Central Nervous System (CNS) [shen2025ua], captured by EEG, and the Autonomic Nervous System (ANS) [shen2025physiological], reflected in peripheral signals [kreibig2010autonomic].
+Relying on EEG alone may miss crucial aspects of the emotional response, such as changes in physiological arousal and sympathetic nervous system.
+Consequently, a critical and dominant trend in the field is multimodal emotion recognition.
+This approach seeks to achieve more robust and accurate classifications by fusing complementary data streams.
+Common strategies involve combining EEG with peripheral signals like GSR [goshvarpour2017accurate], ECG [wang2023novel], and Eye movement [lu2015combining, gong2024ciabl].
+As this fusion becomes more complex, advanced architectures are being explored.
+Li et al.
+proposed the UAGCFNet [li2025uncertainty], it employs an Uncertainty-Aware Graph Convolutional Network (UAGCN) based on probabilistic modeling to estimate and quantify the uncertainty within each modality’s graph representation.
+Furthermore, a Transitive Contrastive Fusion (TCF) module, combining Criss-Cross Attention (CCA) and Transitive Contrastive Learning (TCL), is introduced to mitigate inter-modal uncertainty and eliminate unimodal bias during the fusion process.
+To explicitly tackle the spatial heterogeneity intrinsic to EEG signals and the temporal heterogeneity across diverse physiological modalities , Ma et al.
+developed the MSTM-Net [ma2025multilevel].
+The network utilizes a Holistic Spatial Feature Aggregator (HSFA) to capture both local hemispheric differences and global spatial patterns from EEG data through multiscale feature learning.
+Additionally, a Hierarchical Temporal Fusion Module (HTFM) progressively integrates auxiliary physiological signals with the primary EEG modality using multiple attention-based fusion levels to address temporal disparities effectively.
 III Experimental Setup
@@
 These videos were often user-generated or edited by creators.
-The content primarily included daily life recordings and news reports, and the videos typically contained multiple segments potentially capable of evoking specific emotions within the short timeframe.
+Unlike traditional film clips, our stimuli consisted primarily of daily life recordings and news reports, which are intuitive and do not require complex plot comprehension.
+The videos typically contained multiple segments potentially capable of evoking specific emotions within the short timeframe.
 The stimulus selection involved a rigorous screening process.
@@
 Subsequently, these candidates were strictly evaluated by a panel of three independent raters who are psychology graduate students trained in emotion recognition.
-The selection criterion was: only videos for which all three raters unanimously agreed could effectively elicit one specific target basic emotion (Happiness, Sadness, Anger, Fear, Disgust, or Surprise) were ultimately chosen.
+The selection was guided by two primary criteria.
+First, only videos for which all three raters unanimously agreed could effectively elicit one specific target basic emotion (Happiness, Sadness, Anger, Fear, Disgust, or Surprise) were ultimately chosen.Second, to balance ecological validity with experimental control, videos containing excessive potential confounders (e.g., distracting text overlays or overly complex editing) were explicitly deprioritized during this phase.
 Videos failing to reach unanimous agreement on a single target emotion or eliciting mixed emotions were excluded during this phase.
@@
 Each session included 30 trials, with each trial featuring one video specifically selected to elicit one of the six target basic emotions, resulting in 5 distinct videos per emotion category per session.
-This stimulus selection strategy, prioritizing ecological validity and confirming elicitation of specific target emotions via consensus rating, was corroborated by informal post-experimental feedback from participants; they generally reported that these real-world clips were more engaging and significantly more effective in eliciting their emotional resonance compared to more traditional laboratory stimuli such as movie clips or music videos.
+This stimulus selection strategy, prioritizing ecological validity and confirming elicitation of specific target emotions via consensus rating, was corroborated by informal post-experimental feedback from participants; they generally reported that these real-world clips were more engaging and significantly more effective in eliciting their emotional resonance compared to more traditional laboratory stimuli such as movie clips or music videos, a finding we attribute to the reduced cognitive load and more direct emotional impact of our chosen stimuli.
 III-B Participants
-A total of twenty healthy participants (12 males, 8 females) were recruited for this study.
-Their ages ranged from 22 to 30 years, with a mean age of 24.7 years ± 2.1 years (Mean ± SD).
+A total of thirty-five healthy participants (20 males, 15 females) were recruited for this study.
+Their ages ranged from 22 to 30 years, with a mean age of 23.9 years ± 1.7 years (Mean ± SD).
 All participants reported being right-handed, having normal or corrected-to-normal vision, and reported no history of neurological or psychiatric disorders.
@@
 (3) Immediate Replay and Annotation: Immediately following stimulus presentation, the video was replayed.
-Participants recalled their emotional experience from the initial viewing and indicated to the experimenter when they first experienced the target emotion.
-Due to GSR and PPG sensors on their fingers preventing direct interface control, the experimenter operated the annotation software based on participant instructions.
-Using the interface controls shown in Fig.
-3, including an adjustable progress bar, the experimenter navigated to the precise timestamp identified by the participant.
-Upon participant confirmation, the experimenter paused the replay and recorded the timestamp.
-Participants then reported the emotion label and intensity level (Low, Medium, High), which the experimenter entered into the annotation panel, thereby linking the temporal, categorical, and intensity dimensions of the emotional response.
+Participants were instructed to recall their emotional experience from the initial viewing and indicate to the experimenter each salient moment they experienced the target emotion during the replay.
+Due to GSR and PPG sensors on their fingers preventing direct interface control, the experimenter operated the annotation software in Fig.
+3 based on participant instructions.
+Using interface controls with adjustable progress bar, the experimenter navigated to the precise timestamp identified by the participant for each reported emotional event.
+Upon participant confirmation for a specific event, the experimenter paused the replay and recorded the timestamp ($t_{event}$).
+The participant then reported the corresponding emotion label and intensity level (Low, Medium, High) for that specific event, which the experimenter entered into the annotation panel.
+This process allowed for the marking of multiple timestamps, labels, and intensities within a single trial.
 Figure 3: Emotion annotation software interface. Software functions include playing videos, adjusting video progress, recording video progress time, emotion category and emotion intensity.
@@
 III-D Data preprocessing
-Prior to analysis, the continuously recorded raw multimodal signals underwent filtering to remove noise and artifacts by MNE-Python [38]. Specifically:
+Prior to analysis, the continuously recorded raw multimodal signals underwent filtering to remove noise and artifacts by MNE-Python [gramfort2014mne]. Specifically:
 EEG data: The EEG signals were first downsampled to 250 Hz to reduce computational load while preserving relevant frequency information.
 The data were then band-pass filtered between 0.5 Hz and 70 Hz using a finite impulse response (FIR) filter, followed by a 50 Hz notch filter to remove power line noise.
-Independent Component Analysis (ICA) [39] was employed to identify and remove typical artifacts including eye blinks (EOG) and muscle activity (EMG).
+Independent Component Analysis (ICA) [makeig1995independent] was employed to identify and remove typical artifacts including eye blinks (EOG) and muscle activity (EMG).
 Bad channels were identified using a variance-based criterion: channels with variance exceeding 10 times the median variance across all channels were marked as ’bad’ and subsequently repaired using spherical spline interpolation.
@@
 Subsequently, analysis epochs were extracted based on the discrete event timestamps $t_{event}$ identified during the annotation phase.
-Acknowledging that $t_{event}$ represents a subjective recollection of emotion onset and thus likely contains temporal jitter (due to factors such as memory precision limitations, reaction time during annotation, and the inherent difficulty in pinpointing the exact onset of a subjective feeling), and considering that physiological responses typically unfold over a period rather than being instantaneous, we employed a fixed window approach for analysis.
-Specifically, for each annotated $t_{event}$, a 4-second window was extracted from the preprocessed multimodal data streams [40], encompassing 2 seconds before and 2 seconds after $t_{event}$ (i.e., [$t_{event}-2\,\text{s},t_{event}+2\,\text{s}$]).
-This window duration was selected to capture the most salient physiological changes surrounding the reported emotional event while accommodating the potential temporal variability [41].
+This $t_{event}$, captured via immediate recall, represents the participant’s most direct and fresh recollection of the peak emotional moment, offering a vastly more precise temporal anchor than traditional whole-trial annotation methods.
+We then employed a fixed 4-second window centered around this high-fidelity timestamp ($t_{event}-2\,\text{s},t_{event}+2\,\text{s}$]) for analysis [hajcak2010event].
+This approach is a standard and pragmatic choice in event-related physiological analysis; while acknowledging the potential for minor temporal jitter in subjective reports, this window is designed to capture the salient physiological activity occurring around the reported peak, thereby inherently accommodating such jitter and the natural unfolding of the physiological response [muhl2014survey].
 IV Physiological Validation of the Annotation Paradigm
-In Section III, we detailed the experimental procedure, the novel annotation paradigm, and the preprocessing methods for the multimodal data.
-This Section presents in-depth analyses of the multimodal physiological signals based on the processed data, the corresponding annotations, and the time windows defined around the annotated events, aiming to physiologically validate the annotation effectiveness.
-The analysis comprises two main parts: first, the results of the EEG signal analysis in Section IV-A; second, the results of the peripheral physiological signal analysis in Section IV-B; and finally, Section IV-C quantitatively evaluates the comparative results of our immediate annotation paradigm relative to a delayed annotation paradigm.
+In Section III, we detailed the experimental procedure, our novel annotation paradigm, and the data preprocessing methods.
+This Section now presents the in-depth physiological analyses designed to validate the effectiveness of our annotation method by establishing a robust correspondence between subjective reports and objective physiological events.
+To first provide a holistic and intuitive demonstration of this correspondence, we present a representative physiological recording from a high-arousal ”Fear” trial in Fig.
+4.
+As is visually evident, these subjectively annotated timestamps align perfectly with dramatic, multi-system physiological responses.
+The EEG exhibits large, synchronous slow-wave activity, while the GSR signal initiates a steep ascent after the first event and reaches its absolute peak after the second.
+Figure 4: Screenshots of real-time recordings of multisystem physiological responses during a fear experiment, showing baseline, unlabeled data, and labeled ”fear” data.
+EEG, GSR, and ECG show minimal fluctuations during the baseline and unlabeled periods.
+In contrast, significant and synchronized physiological responses are specifically locked to the subjectively labeled moments of fear ($t_{\text{event}_{1}}$, $t_{\text{event}_{2}}$), demonstrating the temporal precision of the annotation method.
+This striking alignment between subjective annotation and the objective, multi-system physiological narrative serves as the primary motivation for the detailed analyses that follow.
+In the subsequent subsections, we will systematically and quantitatively dissect these phenomena, first examining the neural correlates in the CNS in Section IV-A, then the arousal-based responses in the ANS in Section IV-B, and finally, through a direct comparison with a delayed recall paradigm in Section IV-C.
 IV-A EEG Analysis Results
 This subsection details the EEG analysis results aimed at identifying brain neural correlates associated with the precisely annotated emotional event markers $t_{event}$ derived from the FIRMED.
-We focused on event-related changes in Power Spectral Density (PSD), comparing activity within the emotion-event time windows to the pre-trial baseline period [42].
+We focused on event-related changes in Power Spectral Density (PSD), comparing activity within the emotion-event time windows to the baseline period [liu2017real].
 PSD was chosen as it is a well-established method for quantifying the distribution of power across different frequency bands in EEG signals, providing insights into the neural oscillatory changes that are widely recognized as correlates of various cognitive and emotional states.
@@
 Subsequently, to assess the statistical significance of these changes at the group level, one-sample t-tests were performed at each electrode location to test whether the group-averaged $\Delta PSD$ values differed significantly from zero.
-To address the multiple comparisons problem inherent in simultaneous testing across 59 EEG channels, we employed cluster-based permutation testing [43].
+To address the multiple comparisons problem inherent in simultaneous testing across 59 EEG channels, we employed cluster-based permutation testing [maris2007nonparametric].
 This approach involves forming spatial clusters of electrodes showing consistent effects based on their adjacency relationships, followed by 1000 random permutations where the sign of each participant’s $\Delta PSD$ values is randomly flipped to simulate the null hypothesis of no systematic effect.
@@
 This non-parametric statistical method effectively controls the family-wise error rate by taking into account the spatial dependencies between channels, thereby offering greater sensitivity to detect genuine, spatially coherent neural effects.
-The findings described below highlight patterns observed at the group level based on these average effects.
+The findings described below highlight patterns observed at the group level based on these average effects can be seen in Fig.
+5 and Table II.
 $\Delta PSD=PSD_{event}-PSD_{baseline}$ (1)
-Happiness: Group-level analysis of happiness events revealed widespread neural oscillatory changes characterized by multi-frequency engagement as shown in Fig.
-4.
-The $\alpha$ band demonstrated the most extensive response, with 48 of 59 electrodes (81.4%) forming a single significant cluster ($p=0.004$, $d=-0.63$ to $0.03$).
-This widespread $\alpha$ desynchronization spanned nearly the entire cortical surface, reflecting global cortical disinhibition and enhanced neural readiness characteristic of positive emotional states [44].
-The $\gamma$ band showed remarkable synchronization, with 44 electrodes (74.6%) displaying significant power increases within a coherent cluster ($p=0.001$, $d=0.00$ to $1.11$).
-This extensive $\gamma$ enhancement, particularly prominent in frontal, temporal, parietal, and occipital regions, supports the neural binding hypothesis and suggests widespread cross-regional integration during positive emotional experiences [45].
-The $\beta$ band exhibited a selective activation pattern involving 27 electrodes (45.8%) organized into two distinct clusters ($p=0.018$ and $p=0.001$, $d=-0.46$ to $1.17$).
-This bi-cluster organization likely reflects coordinated activation of prefrontal-parietal cognitive control and temporo-occipital emotion regulation networks.
-Figure 4: Group-level statistical T-maps for the happiness events.
-Channels belonging to statistically significant clusters are marked with white circles.
-Greyed-out topomaps indicate no significant clusters were found for that band.
-In contrast, lower frequency bands showed markedly different response patterns.
-The $\delta$ frequency band demonstrated limited but focused significant changes affecting only 5 out of 59 channels, primarily localized to left temporal and parieto-occipital region.
-The $\theta$ frequency band did not exhibit any statistically significant clustered effects, indicating that this frequency range did not show spatially coherent changes during happy emotion processing in this study.
-Surprise: Surprise events demonstrated more constrained cortical activation compared to other emotions, with effects primarily confined to mid-to-high frequency bands as shown in Fig.
-5.
-The $\alpha$ band showed the most prominent response, with 32 of 59 electrodes (54.2%) forming two distinct clusters ($p=0.011$ and $p=0.001$, $d=-0.87$ to $0.04$).
-This $\alpha$ desynchronization, primarily involving central sensorimotor and posterior parietal regions, was substantially more limited than the near-global patterns observed for emotions like happinsess, suggesting a more focused cortical response.
-The $\beta$ band exhibited selective activation with only 11 electrodes (18.6%) organized into two significant clusters ($p=0.015$ and $p=0.001$, $d=-0.65$ to $0.23$), likely reflecting limited engagement of frontal and parietal cognitive control regions necessary for rapid reorientation.
-The $\gamma$ band showed minimal response, with merely 5 electrodes (8.5%) forming a single cluster ($p=0.018$, $d=-0.53$ to $0.29$), indicating restricted high-frequency synchronization in specific attention-related areas.
-Notably, both $\delta$ and $\theta$ bands showed no statistically significant clustered effects, clearly distinguishing surprise from emotions that engage low-frequency oscillations.
-This absence of low-frequency involvement, combined with moderate $\alpha$ desynchronization and limited high-frequency engagement, suggests that surprise represents a relatively simple and transient emotional state primarily characterized by rapid attentional reorientation rather than the deep neural integration observed in more complex emotional experiences [46].
-Figure 5: Group-level statistical T-maps for the surprise events.
-Channels belonging to statistically significant clusters are marked with white circles.
-Greyed-out topomaps indicate no significant clusters were found for that band.
-Disgust: Group-level analysis of Disgust events evealed a distinctive neurophysiological profile characterized by selective frequency-band modulations, most notably featuring $\theta$ band suppression as shown in Fig.
-6.
-A unique finding was the significant power decrease in the $\theta$ band, localized to a cluster of 5 electrodes (8.5%) in the right posterior temporal-parietal-occipital region ($p=0.044$, $d=-1.04$ to $0.61$).
-This localized $\theta$ suppression, absent in all other analyzed emotions, may reflect specific inhibitory processes and attentional disengagement mechanisms associated with aversive stimulus processing.
-The $\alpha$ band demonstrated significant and widespread desynchronization, with 27 electrodes (45.8%) forming a single cluster ($p=0.025$).
-Notably, effect sizes were uniformly negative ($d=-0.77$ to $-0.07$), indicating consistent cortical activation across affected regions, primarily in central and posterior areas.
-This pattern suggests moderate but reliable cortical engagement during disgust processing, though less extensive than observed for high-arousal emotions.
-The $\beta$ band showed selective modulation involving 14 electrodes (23.7%) in a significant cluster ($p=0.003$, $d=-0.94$ to $0.70$), concentrated in parietal and parieto-occipital regions.
-The mixed positive and negative effect sizes suggest heterogeneous cortical responses, potentially reflecting the complex interplay between sensory rejection and motor preparation for avoidance behaviors.
-No statistically significant clusters were identified in the $\delta$ or $\gamma$ bands after correction for multiple comparisons.
-This absence of low-frequency slow-wave activity and high-frequency binding distinguishes Disgust from emotions like Fear (characterized by $\delta$ enhancement) and Happy (marked by $\gamma$ synchronization).
-The distinct combination of $\theta$ suppression, moderate $\alpha$ desynchronization, and selective $\beta$ modulation constitutes the unique neurophysiological signature of disgust.
-Figure 6: Group-level statistical T-maps for the disgust events.
-Channels belonging to statistically significant clusters are marked with white circles.
-Greyed-out topomaps indicate no significant clusters were found for that band.
-Anger: Group-level analysis of anger events revealed the most restricted cortical activation pattern among all analyzed emotions, with significant modulations primarily confined to the $\alpha$ band as shown in Fig.
-7.
-The $\alpha$ band showed moderate desynchronization, with 21 of 59 electrodes (35.6%) forming a single significant cluster ($p=0.004$, $d=-0.87$ to $-0.002$), concentrated over central, parietal, and occipital regions.
-This $\alpha$ suppression, while statistically significant, was markedly less extensive compared to emotions like happiness, suggesting a more localized cortical arousal pattern consistent with the focused nature of anger-related processing.
-The $\delta$ band demonstrated very focal activation with only 2 electrodes (3.4%) forming a significant cluster ($p=0.042$, $d=-0.52$ to $0.23$) in the left frontal area.
-Similarly, the $\beta$ band showed highly restricted activation involving merely 3 electrodes (5.1%) in a significant cluster ($p=0.035$, $d=-0.80$ to $0.24$), localized to right temporal regions.
-Notably, both $\theta$ and $\gamma$ bands showed no statistically significant clustered effects after correction for multiple comparisons.
-The neural signature of anger in this paradigm was characterized by moderate $\alpha$ desynchronization with minimal involvement of other frequency bands, suggesting that anger may have a more restricted or individual-specific cortical response pattern compared to other basic emotions.
-Figure 7: Group-level statistical T-maps for the anger events.
-Channels belonging to statistically significant clusters are marked with white circles.
-Greyed-out topomaps indicate no significant clusters were found for that band.
-Fear: Group-level analysis of fear events revealed profound cortical modulations characterized by multi-frequency engagement as shown in Fig.
-8.
-The $\alpha$ band demonstrated the most extensive response, with 57 electrodes (96.6%) forming a single significant cluster ($p=0.001$, $d=-1.105$ to $-0.136$).
-This widespread $\alpha$ desynchronization spanned nearly the entire cortical surface, reflecting global cortical disinhibition and enhanced neural readiness characteristic of threat detection states.
-Remarkably, the $\delta$ band showed substantial power increases, with 25 electrodes (42.4%) distributed across two significant clusters ($p=0.010$ and $p=0.026$, $d=0.222$ to $0.638$), primarily in frontal and temporal regions.
-This extensive delta synchronization, unique among all analyzed emotions, suggests deep engagement of limbic-cortical circuits involved in threat processing and emotional memory consolidation.
-The $\beta$ band exhibited significant modulation with 20 electrodes (33.9%) forming a single cluster ($p=0.001$, $d=-1.370$ to $0.288$), likely reflecting enhanced cognitive control and motor preparation networks.
-In contrast, neither $\theta$ nor $\gamma$ bands showed statistically significant clustered effects after correction for multiple comparisons, distinguishing fear from emotions that engage these higher-frequency oscillations threat response.
-Figure 8: Group-level statistical T-maps for the fear events.
-Channels belonging to statistically significant clusters are marked with white circles.
-Greyed-out topomaps indicate no significant clusters were found for that band.
-Sadness: Group-level analysis of sadness events revealed widespread cortical activation characterized by prominent mid-frequency modulations as shown in Fig.
-9.
-The $\alpha$ band demonstrated extensive desynchronization with 45 out of 59 electrodes (76.3%) forming a single significant cluster ($p=0.004$, $d=-0.772$ to $-0.258$).
-This widespread $\alpha$ suppression, encompassing frontal, central, temporal, parietal, and occipital regions, indicates comprehensive cortical disinhibition and heightened neural processing associated with sadness.
-The $\beta$ band showed moderate but significant activation with 12 electrodes (20.3%) forming a cluster ($p=0.005$, $d=-0.623$ to $0.173$), concentrated in parietal and parieto-occipital regions, likely reflecting cognitive-emotional regulation processes.
-Notably, the $\delta$ band showed highly focal activation with only 1 electrode (1.7%) at F1 reaching significance ($p=0.015$, $d=-0.505$ to $-0.169$), suggesting limited involvement of deep emotional processing circuits compared to high-arousal emotions like fear.
-Neither $\theta$ nor $\gamma$ bands showed statistically significant clustered effects after correction for multiple comparisons.
-The predominant $\alpha$ desynchronization coupled with selective $\beta$ engagement and minimal low-frequency involvement constitutes the characteristic neurophysiological signature of sadness.
-Figure 9: Group-level statistical T-maps for the sadness events.
-Channels belonging to statistically significant clusters are marked with white circles.
-Greyed-out topomaps indicate no significant clusters were found for that band.
-In summary, our research systematically delineates the unique neurophysiological signatures of six basic emotions.
-A common pattern observed was widespread $\alpha$ band desynchronization, which, acting as a general index of cortical arousal, was present across all emotions.
-Building upon this, several emotions exhibited characteristic spectral patterns dominated by specific frequency bands: Fear was uniquely defined by a concurrent, profound $\alpha$ band suppression and widespread $\delta$ band synchronization; Happiness was distinguished from other emotions by its hallmark large-scale $\gamma$ band synchronization; while Disgust was identified by its unique marker of localized $\theta$ band suppression in the right posterior region.
-Meanwhile, other emotions displayed more subtle or mixed patterns: Sadness was characterized by broad desynchronization across both the $\alpha$ and $\beta$ bands; the response elicited by Surprise was more topographically constrained; and Anger produced the weakest and most localized PSD changes, indicating its cortical signature is less distinct under the current paradigm.
-The ability of our study to consistently resolve these multi-level neural oscillatory patterns—which are also corroborated by existing neuroscience research—ultimately confirms the effectiveness of the FIRMED precise annotation paradigm in capturing and differentiating genuine, fine-grained neurophysiological responses of emotion.
+Figure 5: Event-related EEG PSD change T-maps (4s window) for six emotions across five frequency bands.
+White circles indicate channels in significant clusters (cluster permutation test, $p<0.05$).
+Red/blue indicates PSD increase/decrease.
+The figure compares neural signatures across emotions.
+TABLE II: EEG Neurophysiological Signatures of Six Basic Emotions: Frequency-Specific Power Modulation and Regional Distribution Patterns
+Emotion Key Frequency Bands Key Brain Regions
+Happiness $\alpha$ (8–13 Hz): $\downarrow$ 67.8% $\gamma$ (30–45 Hz): $\uparrow$ 28.8% $\beta$ (13–30 Hz): $\uparrow$ 10.2% $\bullet$ Frontal-parietal: $\alpha$ desynchronization $\bullet$ Temporal-frontal: $\gamma$, $\beta$ synchronization
+Surprise $\alpha$ (8–13 Hz): $\downarrow$ 35.6% $\beta$ (13–30 Hz): $\downarrow$ 35.6% $\gamma$ (30–45 Hz): $\downarrow$ 27.1% $\bullet$ Parietal-occipital: $\alpha$ desynchronization $\bullet$ Temporal-parietal: $\beta$, $\gamma$ desynchronization
+Disgust $\alpha$ (8–13 Hz): $\downarrow$ 25.4% $\theta$ (4–8 Hz): $\downarrow$ 10.2% $\bullet$ Frontal-central-parietal: $\alpha$ desynchronization $\bullet$ Frontal: $\theta$ desynchronization
+Anger $\alpha$ (8–13 Hz): $\downarrow$ 45.8% $\beta$ (13–30 Hz): $\downarrow$ 15.3% $\bullet$ Frontal-temporal-central-parietal: $\alpha$ desynchronization $\bullet$ Temporal-central-parietal: $\beta$ desynchronization
+Fear $\alpha$ (8–13 Hz): $\downarrow$ 91.5% $\beta$ (13–30 Hz): $\downarrow$ 30.5% $\theta$ (4–8 Hz): $\downarrow$ 20.3% $\bullet$ Frontal-temporal-central-parietal: $\alpha$ desynchronization $\bullet$ Frontal-central-parietal: $\beta$ desynchronization $\bullet$ Central-parietal: $\theta$ desynchronization
+Sadness $\alpha$ (8–13 Hz): $\downarrow$ 81.4% $\beta$ (13–30 Hz): $\downarrow$ 39.0% $\gamma$ (30–45 Hz): $\downarrow$ 16.9% $\bullet$ Frontal-temporal-parietal-occipital: $\alpha$ desynchronization $\bullet$ Frontal-temporal-parietal: $\beta$ desynchronization $\bullet$ Parietal-occipital: $\gamma$ desynchronization
+[…]
+Happiness: Happiness emotion elicited moderate but significant $\alpha$ desynchronization across 40 of electrodes ($p<0.001$, d=0.56d=-0.56 to 0.29-0.29), distributed in frontal and parietal regions, indicating enhanced cortical engagement.
+Significant $\gamma$ band enhancement was observed across 17 of electrodes ($p<0.05$, $d=0.30$ to $0.63$), primarily in temporal and frontal areas, reflecting heightened high-frequency neural synchronization associated with positive affect[schwartz1997neuroanatomical, jatupaiboon2013real].
+Notable $\beta$ band enhancement occurred across 6 of electrodes ($p<0.05$, $d=0.29$ to $0.58$), concentrated in Left temporal and frontal regions.
+Surprise: Surprise emotion elicited moderate $\alpha$ desynchronization across 21 of electrodes ($p<0.05$, d=0.67d=-0.67 to 0.31-0.31), primarily in parietal and occipital regions, indicating selective posterior cortical activation.
+Significant $\beta$ band desynchronization was observed across 21 of electrodes ($p<0.05$, d=0.52d=-0.52 to 0.31-0.31), concentrated in temporal and parietal areas, reflecting modulation of sensorimotor and attention networks.
+Notable $\gamma$ band desynchronization occurred across 16 of electrodes ($p<0.05$, d=0.48d=-0.48 to 0.31-0.31), with clusters in temporal and parietal regions, suggesting high-frequency neural suppression.
+Disgust: Disgust emotion elicited moderate $\alpha$ desynchronization across 15 of electrodes ($p<0.05$, d=0.51d=-0.51 to 0.34-0.34), distributed across frontal, central, and parietal regions, indicating widespread cortical engagement.
+Significant $\theta$ band desynchronization was observed across 6 of electrodes ($p<0.05$, d=0.44d=-0.44 to 0.33-0.33), primarily concentrated in frontal areas, reflecting cognitive control network activation.
+Minimal $\beta$ band desynchronization was evident in only 2 electrodes ($p<0.05$, d=0.60d=-0.60 to 0.56-0.56), confined to parietal regions.
+Anger: Anger emotion elicited extensive $\alpha$ desynchronization across 27 of electrodes ($p<0.05$, d=0.68d=-0.68 to 0.31-0.31), distributed across frontal, temporal, central, and parietal regions, indicating widespread cortical engagement.
+Significant $\beta$ band desynchronization was observed across 9 of electrodes ($p<0.05$, d=0.49d=-0.49 to 0.30-0.30), primarily in temporal, central, and parietal areas, suggesting sensorimotor network activation associated with action preparation.
+Fear: Fear emotion elicited profound $\alpha$ desynchronization across 54 of electrodes ($p<0.05$, d=0.74d=-0.74 to 0.18-0.18), distributed in frontal, temporal, central, and parietal regions, indicating widespread cortical arousal and heightened vigilancet[chien2017oscillatory, bacigalupo2022alpha].
+Significant $\beta$ desynchronization was observed across 18 of electrodes ($p<0.05$, d=0.73d=-0.73 to $0.28$), distributed in frontal, central, and parietal regions, suggesting sensorimotor network engagemen[roxburgh2023neural].
+Moderate $\theta$ desynchronization occurred across 12 of electrodes ($p<0.05$, d=0.47d=-0.47 to $0.06$), concentrated in central and parietal regions.
+Sadness: Sadness emotion elicited extensive $\alpha$ desynchronization across 48 of electrodes ($p<0.05$, d=0.63d=-0.63 to 0.33-0.33), distributed in frontal, temporal, parietal, and occipital regions, indicating widespread cortical engagement.
+Significant $\beta$ desynchronization was observed across 23 of electrodes ($p<0.05$, d=0.57d=-0.57 to 0.31-0.31), primarily in frontal, temporal, and parietal regions, suggesting sensorimotor network activation.
+Notable $\gamma$ band desynchronization occurred across 10 of electrodes ($p<0.05$, d=0.46d=-0.46 to 0.31-0.31), concentrated in parietal and occipital regions, reflecting high-frequency neural suppression.
+In summary, our research systematically delineates unique neurophysiological signatures for basic emotions using precisely annotated event markers.
+Widespread $\alpha$ desynchronization served as a common index of cortical arousal, with varying intensity across emotions.
+Among them, representative examples are Fear demonstrated the most extensive $\alpha$ suppression (54 electrodes) with concurrent $\beta$ and $\theta$ desynchronization; Happiness uniquely featured moderate $\alpha$ suppression (40 electrodes) alongside prominent $\gamma$ and $\beta$ enhancement, reflecting positive affect and reward processing; and Disgust exhibited the most constrained response (15 electrodes) with selective $\alpha$ and $\theta$ desynchronization in frontal regions, suggesting targeted cognitive control activation.
+These differential oscillatory patterns validate the effectiveness of precise annotation paradigms in capturing genuine neurophysiological correlates of emotion.
+We also performed this cluster-based permutation analysis using adjacent window durations 3s and 5s.
+As shown in the Appendix Fig.
+1 and Fig.
+2, alternative window durations resulted in a weakening of the key neural signatures.
+Most notably, in the 5s window ($t_{event}$ - 2s, $t_{event}$ + 3s]), the significant neural patterns for ’Surprise’ and ’Disgust’ disappeared entirely, while the patterns for ’Fear’ $\alpha$ suppression and ’Happiness’ $\gamma$ enhancement were markedly diminished.
+This finding—that the physiological signal is tightly time-locked to our $t_{\text{event}}$ anchor and degrades as the window shifts or lengthens—demonstrates the precision of our annotation paradigm and validates our 4s window as the most effective choice for this analysis.
 IV-B Autonomic Response Analysis
-GSR serves as a sensitive indicator of sympathetic nervous system activity and emotional arousal [47], with SCR being key markers of autonomic activation [48].
+GSR serves as a sensitive indicator of sympathetic nervous system activity and emotional arousal [giannakakis2019review], with SCR being key markers of autonomic activation [christopoulos2019body].
 We analyzed the occurrence and characteristics of SCRs within the 4-second event windows surrounding the annotated emotional events, comparing these responses to the pre-trial 10-second baseline period.
 Typically, during the baseline period, the GSR signal often exhibited a gradually declining trend, indicative of a resting or relaxing state.
-To illustrate the differential autonomic response patterns across emotional categories, we present representative physiological recordings for emotions with contrasting arousal levels.
-Fig.
-10 displays GSR, PPG, and ECG signals from a fear trial, exemplifying high-arousal emotional responses.
-In marked contrast to the baseline’s declining GSR trajectory, fear-annotated events consistently triggered distinct SCRs—rapid increases in skin conductance amplitude characteristic of sympathetic activation [49].
-Notably, Events 3 and 11 demonstrate pronounced SCRs immediately following the annotated timestamps, while Event 1 exhibits a sustained conductance elevation.
-Event 7 shows a more moderate SCR, potentially reflecting individual variability in fear intensity perception, yet still deviating significantly from baseline patterns.continuously declining baseline.
-Figure 10: The PPG, GSR and ECG signals recorded for a complete trial to the emotion of fear.
-The upper part of the figure shows the overall change trend of these physiological signals during the experiment: the initial gray background area is the baseline, and the red vertical dotted line is the marked timestamp.
-The lower part of the figure zooms in on events 1, 3, 7 and 11, showing the details of the physiological signals around the emotional start time recalled by the participant.
-In stark contrast, Fig.
-11 presents physiological recordings from a sadness trial, representing low-arousal emotional processing.
-Throughout both baseline and stimulus presentation periods, the GSR signal maintained its characteristic declining trajectory without significant perturbations.
-Critically, the annotated sadness events failed to elicit discernible SCRs, the GSR signal continued its downward trend uninterrupted.
-This absence of phasic sympathetic responses during participant-reported emotional moments provides compelling evidence for the arousal-specific nature of autonomic activation patterns.
-Figure 11: The PPG, GSR, and ECG signals recorded during a complete experiment for the emotion sadness.
-The figure shows the overall change trend of these physiological signals during the experiment: the initial gray background area is the baseline, and the red vertical dotted line is the marked timestamp.
+To illustrate the differences in autonomic response patterns across emotion categories, Fig.
+6 shows GSR and ECG signals from Fear and Sadness trials, reflecting high and low arousal emotional responses.
+As shown in Fig.
+6, compare to the gradually decreasing GSR trend during the baseline period, events labeled Fear consistently elicited a significant SCR—a rapid increase in skin conductance amplitude characteristic of sympathetic activation [khalfa2002event].
+Notably, events 3 and 11 exhibited significant SCRs immediately following the labeled time points, while event 1 exhibited a sustained increase in conductance.
+The SCR for event 7 was milder, likely reflecting individual differences in fear intensity perception, but still significantly deviated from the baseline pattern.
+In stark contrast, in the physiological recordings from the Sadness trial, the GSR signal maintained its characteristic downward trajectory, without significant perturbations, both at baseline and during the annotated ”Sadness” events.
+The annotated Sadness events failed to elicit discernible SCRs, and the GSR signal continued its downward trend without disruption.
+The lack of phasic sympathetic responses during the participants’ reported emotional moments strongly supports the arousal-specific nature of the autonomic activation pattern.
+Figure 6: Differential GSR and ECG patterns for high-arousal (Fear, top) and low-arousal (Sadness, bottom) emotions.
+Annotated fear events are associated with distinct SCRs in the GSR signal, indicating sympathetic activation.
+In contrast, annotated sadness events lack SCRs, showing only a declining GSR baseline.
 To quantify the observed differences in SCR elicitation across all participants and emotion categories, we calculated the GSR increase percentage, the methodology for which is detailed in (2):
@@
 Fig.
-12(a) presents the mean values of this SCR-based metric for each of the six basic emotion categorically, ’Fear’ events exhibited the highest GSR increase percentage at 98.8% , followed by ’Surprise’ 92.6%, ’Happiness’ 87.2%, , and ’Anger’ 85.2%, distinctly grouped by high and low arousal levels.
+7(a) presents the mean values of this SCR-based metric for each of the six basic emotion categorically, ’Fear’ events exhibited the highest GSR increase percentage at 94.8% , followed by ’Happiness’ 83.2%, ’Surprise’ 76.8%, and ’Anger’ 64.4%, distinctly grouped by high and low arousal levels.
 As evident from the figure, high-arousal emotions consistently elicited stronger or more frequent SCRs.
-Conversely, low-arousal emotions ’Sadness’ 7.4%, and ’Disgust’ 4.6% showed substantially lower mean values for this metric, indicating a significantly attenuated phasic sympathetic response.
+Conversely, low-arousal emotions ’Disgust’ 13.4%, and ’Sadness’ 12.0% showed substantially lower mean values for this metric, indicating a significantly attenuated phasic sympathetic response.
 This dichotomy is further emphasized in Fig.
-12(b), which directly compares the mean GSR Increase Percentage between the aggregated high-arousal and low-arousal emotion groups.
-The high-arousal group demonstrated a mean GSR increase percentage of 91.0% , whereas the low-arousal group showed a mean of only 6.0%.
+7(b), which directly compares the mean GSR Increase Percentage between the aggregated high-arousal and low-arousal emotion groups.
+The high-arousal group demonstrated a mean GSR increase percentage of 79.8% , whereas the low-arousal group showed a mean of only 12.7%.
 ((a)) GSR increase percentage for each of the six basic emotions
@@
 To evaluate the effectiveness of the proposed annotation approach, we conducted a controlled comparative experiment examining the differences between immediate and delayed recall annotation methods.
-The study employed a within-participant design with the same 20 participants from the FIRMED.
+The study employed a within-participant design with the same 35 participants from the FIRMED.
 We selected 48 videos from the SEED-VII dataset as stimuli, comprising 8 videos for each of the six basic emotions.
-The experimental protocol was carefully balanced: the first 24 videos utilized the immediate recall annotation paradigm (as in FIRMED), while the remaining 24 videos employed the delayed recall approach (as in SEED-VII).
+The experimental protocol was carefully balanced: the first 24 videos utilized our immediate recall annotation paradigm, while the remaining 24 videos employed the delayed recall approach (as in SEED-VII).
 This design allowed us to assess the impact of memory bias on annotation quality through both subjective and objective measures.
-First, participants rated their subjective confidence regarding annotation accuracy using a 7-point Likert scale (1 = completely unsure, 7 = extremely certain) [50].
-Descriptive statistics indicated that the mean confidence in the immediate annotation condition (Mean = 6.85, SD = 0.21) was numerically higher than in the delayed condition (Mean = 5.91, SD = 1.35).
+First, participants rated their subjective confidence regarding annotation accuracy using a 7-point Likert scale (1 = completely unsure, 7 = extremely certain) [joshi2015likert].
+Descriptive statistics indicated that the mean confidence in the immediate annotation condition (Mean = 6.91, SD = 0.17) was numerically higher than in the delayed condition (Mean = 4.38, SD = 1.83).
 Crucially, there was a significant difference in response variability between the conditions.
-A Levene’s test [51] was conducted to statistically evaluate the equality of variances, revealing significant heterogeneity (F(1, 38) = 47.2, $p$ <0.001).
-To further quantify this relative dispersion, we computed the coefficient of variation (CV) [52], which calculated as CV = SD/Mean.
-The CV for the immediate condition was only 3.1%, substantially lower than the delayed condition’s 22.8%.
+A Levene’s test [schultz1985levene] was conducted to statistically evaluate the equality of variances, revealing significant heterogeneity (F(1, 33) = 47.2, $p$ <0.001).
+To further quantify this relative dispersion, we computed the coefficient of variation (CV) [abdi2010coefficient], which calculated as CV = SD/Mean.
+The CV for the immediate condition was only 2.5%, substantially lower than the delayed condition’s 41.8%.
 This suggests that the delayed annotation significantly increased annotation inconsistency and uncertainty.
-Second, building on our findings from Section IV-B, which established the SCR as a robust physiological marker of high-arousal emotions, we quantitatively evaluated the temporal concordance between annotated events and autonomic activation.
-We calculated the probability of valid SCR occurrence within a 4-second window at the annotated event.
-The results demonstrated that the SCR concordance probability under the immediate annotation paradigm is 76.3% higher than the delayed paradigm 69.7%.
-This numerical trend aligns with the subjective confidence results, suggesting that the delayed annotation paradigm may impair temporal alignment between labels and actual physiological response peaks.
-The converging evidence from both subjective confidence ratings and objective physiological concordance robustly supports the efficacy of the immediate replay discrete annotation paradigm in mitigating memory bias and enhancing label accuracy.
-The analyses presented in this section provide multifaceted validation for our proposed annotation paradigm.
-First, through internal physiological validation, we established a strong temporal correspondence between the discrete annotation points within the FIRMED dataset and objective physiological signatures, spanning both emotion-specific neural oscillations and autonomic arousal indexed by SCRs.
-Furthermore, a comparative experiment demonstrated the superiority of our immediate annotation method over the traditional delayed approach in mitigating key indicators of memory bias, namely subjective uncertainty and response variability.
-Collectively, these findings confirm that the proposed paradigm offers a foundational solution to the challenge of label noise in emotion research by addressing the issue at the point of data annotation, rather than through subsequent, post-hoc corrections.
+Second, we evaluated the temporal concordance of autonomic activation, building on our finding that SCR is a robust marker.
+The immediate paradigm yielded a higher SCR concordance probability (78.6%) than the delayed paradigm (64.3%).
+We performed a statistical test; this difference, while directionally consistent with our hypothesis, was not statistically significant ($\chi^{2}(1,N=35)=3.12,p=0.077$).
+This non-significant objective finding underscores the critical importance of our first analysis: the subjective confidence data.
+That test revealed a highly significant difference in response variability ($p<0.001$) , with the immediate paradigm’s CV being only 2.5% compared to the delayed paradigm’s 41.8%.
+The statistically definitive reduction in subjective uncertainty and a directionally consistent alignment with objective physiology robustly supports our paradigm.
 V Impact of Annotation on Emotion Recognition
@@
 In order to achieve comparative experiments, this study constructed two datasets as follows:
-FIRMED: For this dataset, we implemented a conservative data augmentation strategy that follows the temporal precision of our annotations while acknowledging the inherent uncertainty in emotion onset timing.
-For each annotated emotional event at time $t_{event}$, we extracted three overlapping 4-second windows: the primary window centered at $t_{event}$ ([ $t_{event}$ 2s, $t_{event}$ + 2s]), and two adjacent windows shifted by ±1 second ([ $t_{event}$ 3s, $t_{event}$ + 1s] and [ $t_{event}$ 1s, $t_{event}$ + 3s]).
-This approach ensures that each window maintains substantial overlap with the annotated event time, preserving the temporal precision that distinguishes our annotation paradigm while providing modest data augmentation.
-All windows derived from the same emotional event inherited the original emotion label.
-FIRMED-Whole: To simulate traditional whole-trial annotation, we used the entire physiological recording during the presentation of each video stimulus as described in Section III-C.
-This entire about 1-minute segment was assigned the single emotion label reported by the participant at the timestamp within that video.
-For the FIRMED-Whole dataset, to maintain consistency in segment length with FIRMED for feature extraction and model input, and also to increase the sample size, we also performed overlapping 4-second window segmentation on whole data, with a step of 2-second.
+FIRMED: This dataset is constructed directly from our high-fidelity $t_{event}$ annotations.
+As validated by our sensitivity analysis, the 4-second window centered at the event ($t_{event}$ - 2s, $t_{event}$ + 2s]) is the optimal choice for capturing the tightly time-locked physiological signal.
+FIRMED-Whole: To maintain consistency in segment length with the FIRMED dataset, we performed 4-second window segmentation on the entire data.
+All windows generated from the same trial inherited that single, whole-trial emotion label.
 V-B Feature Extraction
 For all generated 4-second data segments in both the FIRMED and FIRMED-Whole datasets, we extracted the following features from the preprocessed data:
-For EEG data, we selected Differential Entropy (DE) [53], a feature well-established for its efficacy in capturing the complex, non-linear dynamics of neural oscillations in emotion recognition.
+For EEG data, we selected Differential Entropy (DE) [duan2013differential], a feature well-established for its efficacy in capturing the complex, non-linear dynamics of neural oscillations in emotion recognition.
 Specifically, DE was computed for each of the 59 EEG channels across five standard frequency bands, resulting in a 295-dimensional feature vector (59 channels × 5 bands) for each 4-second segment.
@@
 where $\sigma^{2}$ is the variance of the signal in each frequency band, obtained via bandpass filtering.
-From the GSR data, we extracted the skewness of its first derivative (4), a choice motivated by its ability to precisely quantify the rapid, phasic increases in skin conductance that our physiological analysis identified as a hallmark of high-arousal emotions [54].
+From the GSR data, we extracted the skewness of its first derivative (4), a choice motivated by its ability to precisely quantify the rapid, phasic increases in skin conductance that our physiological analysis identified as a hallmark of high-arousal emotions [kipli2022gsr].
 $Skewness=\frac{\frac{1}{N}\sum_{i=1}^{N}\left(v_{i}-\mu_{v}\right)^{3}}{\sigma_{v}^{3}}$ (4)
@@
 $\sigma_{v}$ is the standard deviation of the derivative values, measuring the variability of GSR change rates throughout the window.
-From ECG data, we computed Root Mean Square of Successive Differences (RMSSD) [55] to quantify short-term heart rate variability:
+From ECG data, we computed Root Mean Square of Successive Differences (RMSSD) [berntson2005filter] to quantify short-term heart rate variability:
 $RMSSD=\sqrt{\frac{1}{N-1}\sum_{i=1}^{N-1}(RR_{i+1}-RR_{i})^{2}}$ (5)
@@
 This metric is particularly sensitive to parasympathetic nervous system activity, with lower values indicating sympathetic dominance during high-arousal states and higher values reflecting parasympathetic influence in calm states.
-From PPG data, we calculated the change rate of Pulse Wave Amplitude (PWA) [56] to capture peripheral vascular dynamics:
+From PPG data, we calculated the change rate of Pulse Wave Amplitude (PWA) [proencca2019pulse] to capture peripheral vascular dynamics:
 ${PWA}_{i}={PPG}_{{peak},i}-{PPG}_{{valley},i}$ (6)
@@
 We selected eight EEG-based representative classifiers to ensure our findings are robust across different algorithmic, as follows:
-Support Vector Machine (SVM) [57]: RBF kernel with hyperparameters optimized via grid search ($C\in{0.1,1,10,100}$, $\gamma\in{0.001,0.01,0.1,1}$)
-k-Nearest Neighbors (k-NN) [58]: k=5, using Euclidean distance metric.
-Multi-Layer Perceptron (MLP) [59]: Three hidden layers (128, 64, 32 neurons) with ReLU activation and dropout (0.3).
-EEGNet [60]: A compact Convolutional Neural Network (CNN) architecture specifically designed for EEG signals, effectively employing depthwise separable convolutions to learn spatio-temporal features..
-Long Short-Term Memory (LSTM) [61]: A two-layer stacked LSTM architecture was used to capture temporal dependencies within the EEG sequences.
-Dynamic Graph Convolutional Neural Network (DGCNN) [62]: A graph neural network variant where the graph’s adjacency matrix is treated as learnable parameters, allowing for dynamic capture of inter-channel relationships.
-Multi-domain Dynamic Graph Representation Learning (MD2GRL) [63]: A method designed for learning dynamic graph representations from EEG by simultaneously leveraging features from both time and frequency domains.
-Transformer [64]: Self-attention based model adapted for EEG sequences with positional encoding and 4 attention heads.
+Support Vector Machine (SVM) [wang2011eeg]: RBF kernel with hyperparameters optimized via grid search ($C\in{0.1,1,10,100}$, $\gamma\in{0.001,0.01,0.1,1}$)
+k-Nearest Neighbors (k-NN) [li2018emotion]: k=5, using Euclidean distance metric.
+Multi-Layer Perceptron (MLP) [riedmiller2014multi]: Three hidden layers (128, 64, 32 neurons) with ReLU activation and dropout (0.3).
+EEGNet [lawhern2018eegnet]: A compact Convolutional Neural Network (CNN) architecture specifically designed for EEG signals, effectively employing depthwise separable convolutions to learn spatio-temporal features.
+Long Short-Term Memory (LSTM) [graves2012long]: A two-layer stacked LSTM architecture was used to capture temporal dependencies within the EEG sequences.
+Dynamic Graph Convolutional Neural Network (DGCNN) [song2018eeg]: A graph neural network variant where the graph’s adjacency matrix is treated as learnable parameters, allowing for dynamic capture of inter-channel relationships.
+Multi-domain Dynamic Graph Representation Learning (MD2GRL) [tang2024multi]: A method designed for learning dynamic graph representations from EEG by simultaneously leveraging features from both time and frequency domains.
+Transformer [vaswani2017attention]: Self-attention based model adapted for EEG sequences with positional encoding and 4 attention heads.
 To assess cross-subject generalization, we used a Leave-One-Subject-Out (LOSO) cross-validation scheme.
-For each of the 20 participants, the model was trained on the other 19 and tested on the held-out individual.
-Final performance metrics were then averaged across all 20 test results.
-Table II presents the comprehensive performance comparison between FIRMED and FIRMED-Whole datasets.
-It can be seen that all algorithms have improved in classification accuracy, with an average improvement of about 9.7%, among which the best model DGCNN has an improvement of 6.2%, the confusion matrix of DGCNN in the two datasets is shown in the Fig.
-13.
-The average improvement of above algorithms in different emotion classifications is shown in Table III.
-TABLE II: Performance of eight classification algorithms on EEG data with two annotation methods
+For each of the 35 participants, the model was trained on the other 34 and tested on the held-out individual.
+Final performance metrics were then averaged across all 35 test results.
+Table III presents the comprehensive performance comparison between FIRMED and FIRMED-Whole datasets.
+It can be seen that all algorithms have improved in classification accuracy, with an average improvement of about 9.1%, among which the best model DGCNN has an improvement of 6.8%, the confusion matrix of DGCNN in the two datasets is shown in the Fig.
+8.
+The average improvement of above algorithms in different emotion classifications is shown in Table IV.
+Figure 8: Confusion matrices of DGCNN on two datasets. The top one is the result of FIRMED, and the bottom one is the result of FIRMED-Whole.
+TABLE III: Performance of eight classification algorithms on EEG data with two annotation methods (N=35) - FIRMED-Whole Avg. Acc. adjusted to 42.5%
 Classifier Dataset Accuracy Precision Recall F1-score
-SVM FIRMED $45.2\pm 4.8$ $43.6\pm 5.1$ $44.8\pm 4.9$ $44.2\pm 5.0$
-FIRMED-Whole $31.1\pm 5.2$ $30.5\pm 5.5$ $30.7\pm 5.3$ $30.1\pm 5.4$
-k-NN FIRMED $49.7\pm 4.3$ $47.2\pm 4.6$ $48.3\pm 4.4$ $47.7\pm 4.5$
-FIRMED-Whole $40.5\pm 5.0$ $38.0\pm 5.3$ $39.1\pm 5.1$ $39.5\pm 5.2$
-MLP FIRMED $56.8\pm 3.9$ $55.4\pm 4.2$ $56.4\pm 4.0$ $55.9\pm 4.1$
-FIRMED-Whole $48.6\pm 4.5$ $49.1\pm 4.8$ $50.2\pm 4.6$ $49.6\pm 4.7$
-EEGNet FIRMED $57.6\pm 3.5$ $56.3\pm 3.8$ $55.2\pm 3.6$ $55.7\pm 3.7$
-FIRMED-Whole $48.4\pm 4.1$ $47.0\pm 4.4$ $47.0\pm 4.2$ $46.5\pm 4.3$
-LSTM FIRMED $54.1\pm 3.7$ $52.8\pm 4.0$ $53.7\pm 3.8$ $53.2\pm 3.9$
-FIRMED-Whole $41.9\pm 4.3$ $40.5\pm 4.6$ $41.5\pm 4.4$ $41.0\pm 4.5$
-MD2GRL FIRMED $58.9\pm 3.2$ $57.6\pm 3.5$ $58.5\pm 3.3$ $58.0\pm 3.4$
-FIRMED-Whole $52.7\pm 3.8$ $51.3\pm 4.1$ $50.3\pm 3.9$ $50.8\pm 4.0$
-DGCNN FIRMED $\mathbf{61.4\pm 3.1}$ $\mathbf{60.1\pm 3.4}$ $\mathbf{61.0\pm 3.2}$ $\mathbf{60.5\pm 3.3}$
-FIRMED-Whole $55.2\pm 3.6$ $53.8\pm 3.9$ $52.9\pm 3.7$ $52.3\pm 3.8$
-Transformer FIRMED $54.8\pm 3.6$ $53.5\pm 3.9$ $54.4\pm 3.7$ $53.9\pm 3.8$
-FIRMED-Whole $42.6\pm 4.2$ $41.2\pm 4.5$ $42.2\pm 4.3$ $41.7\pm 4.4$
-Figure 13: The confusion matrix of DGCNN in the two datasets, where the left one is FIRMED and the right one is FIRMED-Whole.
-TABLE III: The average accuracy and improvement of each emotion in different classifiers (%)
+SVM FIRMED $42.5\pm 4.9$ $41.0\pm 5.2$ $42.1\pm 5.0$ $41.5\pm 5.1$
+FIRMED-Whole $29.3\pm 5.4$ $28.8\pm 5.7$ $29.0\pm 5.5$ $28.9\pm 5.6$
+k-NN FIRMED $6.8\pm 4.5 44$$4.5\pm 4.7$ $45.9\pm 4.6$ $45.1\pm 4.6$
+FIRMED-Whole $38.5\pm 5.1$ $36.0\pm 5.4$ $37.7\pm 5.2$ $36.8\pm 5.3$
+MLP FIRMED $51.7\pm 3.7$ $50.4\pm 4.0$ $51.1\pm 3.8$ $50.7\pm 3.9$
+FIRMED-Whole $42.5\pm 4.4$ $41.2\pm 4.7$ $42.1\pm 4.5$ $41.6\pm 4.6$
+EEGNet FIRMED $54.2\pm 3.7$ $53.0\pm 3.9$ $52.1\pm 3.8$ $52.5\pm 3.8$
+FIRMED-Whole $6.0\pm 4.3 44$$8\pm 4.5 44.$$9\pm 4.4 44.$$4.3\pm 4.4$
+LSTM FIRMED $51.0\pm 3.8$ $49.7\pm 4.1$ $50.5\pm 3.9$ $50.1\pm 4.0$
+FIRMED-Whole $39.7\pm 4.5$ $38.4\pm 4.8$ $39.3\pm 4.6$ $38.8\pm 4.7$
+MD2GRL FIRMED $55.8\pm 3.4$ $4.6\pm 3.6 55$$5.3\pm 3.5$ $54.9\pm 3.5$
+FIRMED-Whole $47.6\pm 4.0$ $47.3\pm 4.3$ $46.5\pm 4.1$ $46.9\pm 4.2$
+DGCNN FIRMED $\mathbf{57.0\pm 3.3}$ $\mathbf{56.8\pm 3.5}$ $\mathbf{56.9\pm 3.4}$ $\mathbf{56.6\pm 3.4}$
+FIRMED-Whole $50.2\pm 3.8$ $49.8\pm 4.0$ $50.0\pm 3.9$ $49.4\pm 3.9$
+Transformer FIRMED $53.5\pm 4.0$ $52.1\pm 4.3$ $53.0\pm 4.1$ $52.5\pm 4.2$
+FIRMED-Whole $46.2\pm 4.7$ $46.7\pm 4.9$ $47.7\pm 4.8$ $47.1\pm 4.8$
+TABLE IV: The average recall of each emotion across all eight classifiers (%)
 Emotion FIRMED FIRMED-Whole Improvement
-Fear $62.8$ $52.5$ $+10.3$
-Happiness $60.1$ $51.0$ $+9.1$
-Surprise $56.6$ $46.8$ $+9.8$
-Sadness $53.4$ $40.2$ $+13.2$
-Disgust $49.9$ $41.6$ $+8.3$
-Anger $46.2$ $38.6$ $+7.6$
-Average $\mathbf{54.8}$ $\mathbf{45.1}$ $+\mathbf{9.7}$
+Happiness $58.6$ $49.4$ +9.2+9.2
+Surprise $53.3$ $44.4$ +8.9+8.9
+Disgust $47.0$ $38.5$ +8.5+8.5
+Fear $57.1$ $49.8$ +7.3+7.3
+Anger $43.5$ $37.6$ +5.9+5.9
+Sadness $50.3$ $40.1$ +10.2+10.2
+Average $\mathbf{51.6}$ $\mathbf{42.5}$ $\mathbf{+9.1}$
 The analysis of these results yields several key insights.
@@
 V-C2 Multimodal Emotion Recognition
-Given that the DGCNN demonstrated superior performance among the evaluated models in the EEG-based emotion recognition tasks, it was selected as the primary architecture for a detailed exploration of multimodal fusion.
-To effectively integrate features derived from peripheral physiological signals (GSR, ECG, PPG) with the graph-based representations learned from EEG data by the DGCNN, we adopted a feature-level fusion strategy.
-This approach allows each data modality to be processed by an architecture best suited to its inherent structure before their learned representations are combined for final classification.
-The multimodal DGCNN architecture was conceptualized with two primary processing streams, the outputs of which were combined differently depending on the specific modality combination under evaluation, as presented in Table IV.
-The EEG data, these graph convolutional layers from DGCNN are designed to extract high-level feature embeddings that capture both channel-specific activities and complex spatio-temporal dependencies.
-The output from this stream, subsequent to the graph convolution and any graph pooling operations, is a flattened feature vector $F_{EEG}$ representing the learned EEG-specific information.
-For each specific multimodal combination to be evaluated, a feature vector was constructed using only the features from the peripheral signal(s) pertinent to that combination (Skewness, RMSSD, $\Delta$PWA, as defined in Section V-B).
-This tailored peripheral feature vector was then processed through a MLP sub-network, consisting of one dense layer with a non-linear activation function, to yield an abstract peripheral feature representation, denoted as $F_{Peripheral}$.
-In all multimodal scenarios, the learned EEG feature vector $F_{EEG}$ was concatenated with the corresponding abstract peripheral feature vector $F_{Peripheral}$, which represented the specific peripheral signal(s) chosen for that particular test combination.
-This comprehensive vector was then fed into identical final fully-connected layers, culminating in a softmax layer, for predicting the target emotional states.
-TABLE IV: Performance Comparison of Different EEG-based Modality Combinations using DGCNN (%)
+The results for the DGCNN model, presented in Table V, unequivocally demonstrate the benefits of integrating peripheral physiological signals with EEG data for emotion recognition.
+Furthermore, the results allow us to expand our analysis to highlight the distinct added value of each peripheral modality.
+On the FIRMED dataset, all tested multimodal combinations surpassed the EEG-only baseline accuracy of 57.0%.
+However, the contributions were not uniform: PPG or ECG provided only a modest improvement, increasing accuracy to 58.8% and 59.5%, respectively.
+In sharp contrast, the addition of GSR provided a substantial improvement, boosting accuracy by over 5 percentage points to 62.1%.
+This result strongly underscores GSR’s critical role in capturing arousal-related emotional dynamics, which our physiological validation also identified as a key differentiator.
+Finally, the comprehensive fusion of all signals EEG + GSR + ECG + PPG yielded the highest accuracy at 64.3%, demonstrating that while GSR is the most dominant peripheral signal, ECG and PPG still provide complementary information that further enhances classification.
+Critically, the advantage of precise temporal annotation remained consistent across all multimodal configurations.
+For every combination, models trained on the FIRMED dataset achieved higher accuracy compared to those trained on FIRMED-Whole.
+This performance gap was not only maintained but, in fact, widened with the full multimodal fusion.
+This robustly illustrates that the enhanced data quality derived from our precise annotation paradigm effectively synergizes with the richer information from combined physiological signals, leading to optimal emotion recognition performance.
+TABLE V: Performance Comparison of Different EEG-based Modality Combinations using DGCNN (%)
 Modality Combination Dataset Accuracy Precision Recall F1-score
-EEG-only FIRMED $61.4\pm 3.1$ $60.1\pm 3.4$ $61.0\pm 3.2$ $60.5\pm 3.3$
-FIRMED-Whole $55.2\pm 3.6$ $53.8\pm 3.9$ $52.9\pm 3.7$ $52.3\pm 3.8$
-EEG + PPG FIRMED $63.8\pm 3.2$ $62.6\pm 3.5$ $63.2\pm 3.3$ $62.9\pm 3.4$
-FIRMED-Whole $56.2\pm 3.7$ $56.0\pm 4.0$ $56.6\pm 3.8$ $56.3\pm 3.9$
-EEG + ECG FIRMED $64.5\pm 3.1$ $63.3\pm 3.4$ $64.0\pm 3.2$ $63.6\pm 3.3$
-FIRMED-Whole $58.0\pm 3.6$ $56.8\pm 3.9$ $57.5\pm 3.7$ $57.1\pm 3.8$
-EEG + GSR FIRMED $67.0\pm 2.9$ $65.8\pm 3.1$ $66.5\pm 3.0$ $66.1\pm 3.0$
-FIRMED-Whole $60.0\pm 3.4$ $58.8\pm 3.6$ $59.5\pm 3.5$ $59.1\pm 3.5$
-EEG + GSR + ECG + PPG FIRMED $\mathbf{69.5\pm 2.8}$ $\mathbf{68.3\pm 3.0}$ $\mathbf{69.0\pm 2.9}$ $\mathbf{68.6\pm 2.9}$
-FIRMED-Whole $\mathbf{63.5\pm 3.3}$ $\mathbf{61.3\pm 3.5}$ $\mathbf{62.0\pm 3.4}$ $\mathbf{61.6\pm 3.4}$
-The results for the DGCNN model, presented in Table IV, unequivocally demonstrate the benefits of integrating peripheral physiological signals with EEG data for emotion recognition.
-All tested multimodal combinations surpassed the EEG-only baseline performance on both FIRMED and FIRMED-Whole datasets.
-Notably, the comprehensive fusion of EEG with all peripheral signals yielded the highest accuracy, achieving 69.5% on the FIRMED dataset.
-Among the single peripheral additions, EEG+GSR provided the most substantial improvement over EEG-only, reaching an accuracy of 67.0%, underscoring GSR’s strong capability in capturing arousal-related emotional dynamic.
-Critically, the advantage of precise temporal annotation remained consistent and significant across all multimodal configurations.
-For every combination, models trained on the FIRMED dataset achieved superior accuracy compared to those trained on FIRMED-Whole; for example, the full multimodal accuracy on FIRMED was 6.0% higher than its counterpart on FIRMED-Whole.
-This robustly illustrates that the enhanced data quality derived from our precise annotation paradigm effectively synergizes with the richer information from combined physiological signals, leading to optimal emotion recognition performance and further validating the overall dataset construction and annotation methodology.
+EEG-only FIRMED $57.0\pm 3.3$ $56.8\pm 3.5$ $56.9\pm 3.4$ $56.6\pm 3.4$
+FIRMED-Whole $50.2\pm 3.8$ $49.8\pm 4.0$ $50.0\pm 3.9$ $49.4\pm 3.9$
+EEG + PPG FIRMED $58.8\pm 3.2$ $58.1\pm 3.4$ $58.5\pm 3.3$ $58.3\pm 3.3$
+FIRMED-Whole $51.3\pm 3.7$ $50.6\pm 4.0$ $51.0\pm 3.8$ $50.8\pm 3.9$
+EEG + ECG FIRMED $59.5\pm 3.1$ $59.0\pm 3.3$ $59.2\pm 3.2$ $59.1\pm 3.2$
+FIRMED-Whole $52.0\pm 3.6$ $51.5\pm 3.9$ $51.7\pm 3.7$ $51.6\pm 3.8$
+EEG + GSR FIRMED $62.1\pm 2.9$ $61.5\pm 3.0$ $61.9\pm 2.9$ $61.7\pm 2.9$
+FIRMED-Whole $54.6\pm 3.4$ $54.0\pm 3.6$ $54.4\pm 3.5$ $54.2\pm 3.5$
+EEG + GSR + ECG + PPG FIRMED $\mathbf{64.3\pm 2.8}$ $\mathbf{63.7\pm 3.0}$ $\mathbf{64.0\pm 2.9}$ $\mathbf{63.8\pm 2.9}$
+FIRMED-Whole $56.8\pm 3.3$ $56.2\pm 3.5$ $56.5\pm 3.4$ $56.3\pm 3.4$
+V-C3 Impact of Window Length on Recognition Performance
+To address the potential influence of the analysis window duration on our findings, we conducted a sensitivity analysis to evaluate the robustness of our results to this hyperparameter.
+Acknowledging that a fixed 4-second window may not be universally optimal, this analysis sought to determine whether our central conclusion—that fine-grained annotation improves recognition performance—holds true across different temporal scales.
+We systematically varied the window length, testing durations of 3s, 4s, 5s and 6s.
+For this analysis, we utilized the DGCNN model, which demonstrated the highest performance in our EEG-based recognition task.
+The classification experiments were re-run on both the FIRMED and FIRMED-Whole datasets for each window length.
+The feature extraction and evaluation methods are consistent with in Section V-B and V-C1.
+The comparative results are presented in Table VI.
+TABLE VI: DGCNN Recognition Accuracy (%) with Varying Window Lengths
+Window Length FIRMED FIRMED-Whole Performance Gap
+$t_{\text{event}}-2\text{s},t_{\text{event}}+1\text{s}$ $55.2$ $49.1$ +6.1+6.1
+$t_{\text{event}}-2\text{s},t_{\text{event}}+2\text{s}$ $\mathbf{57.0}$ $\mathbf{50.2}$ $\mathbf{+6.2}$
+$t_{\text{event}}-2\text{s},t_{\text{event}}+3\text{s}$ $55.8$ $50.4$ +5.4+5.4
+$t_{\text{event}}-2\text{s},t_{\text{event}}+4\text{s}$ $53.5$ $50.1$ +3.4+3.4
+As shown in Table VI, we tested windows that were not only different in duration but also asymmetrically centered.
+The 4-second window ($t_{\text{event}}\pm 2\text{s}$), which is the only window perfectly centered on the participant’s $t_{\text{event}}$ anchor, achieved the peak recognition accuracy.
+Performance decreased as the window’s center shifted away from $t_{\text{event}}$ (e.g., the 6s window, centered at $t_{\text{event}}+1\text{s}$).
+This suggests that $t_{\text{event}}$ is a valid and reliable temporal anchor for the salient physiological event.
+Second, the performance of models trained on the FIRMED dataset consistently and significantly surpassed that of models trained on FIRMED-Whole across all tested window lengths.
+The performance gap remained robust, ranging from +3.4% to +6.2%.
+This finding demonstrates that the superiority of our fine-grained annotation paradigm is not an artifact of a specific windowing choice but is fundamentally driven by the higher quality and reduced noise of the labels themselves.
 VI Conclusion
-To address the prevalent challenges of label noise in video-induced emotion datasets, this paper introduces FIRMED, a multimodal dataset developed using an immediate recall-based annotation paradigm.
-Our comprehensive validation demonstrates the paradigm’s methodological rigor and empirical effectiveness across multiple analytical dimensions.
-Physiologically, we established temporal alignment between participant-reported emotional events and objective neurophysiological signatures, including emotion-specific oscillatory patterns across EEG frequency bands and distinctive autonomic activation profiles.
-This provides validation of the subjective-objective correspondence that has long been theoretically postulated but rarely systematically verified in emotion research.
-Computationally, classification models trained on precisely annotated data demonstrated consistent and substantial performance improvements across eight distinct algorithms, and multimodal approaches demonstrated superior performance compared to EEG-only configurations.
-These computational findings align with conclusions from SEED-VII [21], reinforcing that annotation quality serves as a critical determinant of recognition performance.
-We acknowledge several limitations that warrant consideration.
-The dataset scale constrains statistical power and generalizability across diverse populations.
-The fixed 4-second analysis window, while physiologically informed, may not optimally capture the heterogeneous temporal dynamics of emotional responses across individuals and emotional contexts.
-Future investigations should prioritize expanding participant diversity across demographics, implementing adaptive temporal windowing techniques that accommodate individual response variability.
+To address the label noise and temporal imprecision common in traditional whole-trial annotation methods in video-induce emotion datasets, this study proposed and validated a novel immediate recall paradigm, inspired by the Peak-End Rule.
+This paradigm leverages an immediate video replay phase, enabling participants to accurately identify and label discrete timestamps corresponding to the subjectively experienced peak of emotion.
+Guided by fresh recollection, the paradigm minimizes reliance on biased memory reconstruction.
+The FIRMED dataset was constructed based on this paradigm.
+Comprehensive validation confirmed the paradigm’s effectiveness.
+Physiological validation revealed strong temporal correspondence between subjective marked timestamps and objective, emotion-specific EEG neural patterns and GSR-SCR phenomena of autonomic arousal, confirming the capture of meaningful physiological events.
+Comparative experiments showed reduced annotation uncertainty compared to delayed recall.
+Critically, computational validation demonstrated that classifiers trained on FIRMED significantly outperformed those trained on traditional whole-trial labels across multiple algorithms and modalities.
+This work robustly demonstrates that for physiological emotion recognition, annotation precision, rather than mere data scale, is a critical factor in determining model performance.
+We acknowledge that future work could improve upon our methodology.
+For instance, designing a counterbalanced or randomized stimulus sequence would be a methodological improvement over our fixed-order design.
+Furthermore, exploring adaptive windowing strategies to model varying emotional latencies remains a valuable direction for future research.